### v2

# Meta-learning-based percussion transcription and $t\bar{a}la$ identification from low-resource audio

Abstract
This study introduces a meta-learning-based approach for low-resource Tabla Stroke Transcription (TST) and $t\bar{a}la$ identification in Hindustani classical music. Using Model-Agnostic Meta-Learning (MAML), we address the challenge of limited annotated datasets, enabling rapid adaptation to new tasks with minimal data. The method is validated across various datasets, including tabla solo and concert recordings, demonstrating robustness in polyphonic audio scenarios. We propose two novel $t\bar{a}la$ identification techniques based on stroke sequences and rhythmic patterns. Additionally, the approach proves effective for Automatic Drum Transcription (ADT), showcasing its flexibility for Indian and Western percussion music. Experimental results show that the proposed method outperforms existing techniques in low-resource settings, significantly contributing to music transcription and studying musical traditions through computational tools.
Index Terms:
Automatic Drum Transcription, Model-Agnostic Meta-Learning, Rhythm, Tabla, $T\bar{a}la$, Tabla Stroke Transcription
I Introduction
Music Information Retrieval (MIR) focuses on extracting meaningful features from music recordings, enabling diverse applications in music education, production, and musicology [1]. Among its many tasks, Automatic Music Transcription (AMT) is particularly challenging, as it involves converting audio recordings into symbolic representations such as scores or MIDI files. AMT supports critical tasks such as melody extraction, chord recognition, beat tracking, and rhythm analysis, forming the foundation for advanced music research and applications [2]. While substantial progress has been made in transcribing melodic instruments like the piano and guitar, the transcription of percussive instruments remains under-explored [1]. Furthermore, research progress in the percussion transcription of Hindustani classical music has been significantly slower compared to Western music [3].
The tabla is a key instrument in Hindustani music, serving as the rhythmic backbone and accompanying singers and instrumentalists. It helps maintain the structure and essence of the chosen raga (melodic framework). Tabla Stroke Transcription (TST) is crucial for analysing the complex rhythmic patterns ($t\bar{a}las$) that form the basis of Hindustani music. Accurate TST also aids in preserving and digitising traditional music and creating learning tools for learners and musicians. Like Automatic Drum Transcription (ADT) in Western music, which focuses on drum sound transcription to understand rhythm structures, organise music collections and develop advanced tools, TST aims to identify and classify rhythmic events. However, unlike ADT, TST presents distinct challenges due to the tabla’s unique physical and timbral characteristics compared to drums [1, 4].
TST and ADT are challenging due to the complexity of audio signals in music tracks, where multiple instruments (polyphony) often mask rhythmic sounds. In Western music, bass guitars may overlap with bass drums, while melodic instruments or vocals can obscure tabla strokes in Hindustani music. Transcription methods address these challenges at different levels of complexity: Tabla/Drum-only Transcription (DTD) focuses on isolated recordings, avoiding polyphony. Tabla/Drum Transcription with Percussion (DTP) includes additional percussive elements, adding complexity. The most challenging, Tabla/Drum Transcription with Melodic Instruments (DTM), handles full audio mixes with all instruments and vocals. Since DTM represents most real-world musical tracks, it is the most critical and complex task in transcription.
Most of the present works in ADT rely on deep learning (DL) models due to their ability to handle complex audio mixtures. Similarly, few studies have been conducted on TST, with recent efforts also focusing on DL models [5, 6]. However, DL models require large, annotated datasets for supervised training, which are scarce. Creating these datasets is challenging because annotations must be precisely aligned with the audio, a labour-intensive process requiring skilled musicians. Additionally, many annotated tracks are copyrighted, limiting their public availability. To address this, synthetic data generated from MIDI tracks has been used to produce large-scale training datasets with accurate labels. Despite their large volume, synthetic datasets often perform less effectively than real-world annotated data, making dataset scarcity a significant challenge for DL-based ADT and TST. To overcome this limitation, this paper introduces a new stroke transcription method based on model-agnostic meta-learning (MAML) [7], which operates effectively with minimal data.
TST provides tabla strokes, and $t\bar{a}la$ is typically identified by stroke sequence. While some studies have explored music sequence matching [8, 9], these approaches are not directly applicable to $t\bar{a}la$ identification. In this study, we propose two new $t\bar{a}la$ identification methods.
The key contributions of this work are summarised as follows:
  1. 1.
    A novel meta-learning-based approach for Tabla Stroke Transcription (TST) is proposed and validated on both tabla solo and complete concert recordings. This method is further validated for Automatic Drum Transcription (ADT) tasks using DTD, DTP, and DTM datasets.
  2. 2.
    Two new $t\bar{a}la$ identification methods based on tabla stroke sequences are introduced.
  3. 3.
    A synthetic dataset is curated from the existing mridangam stroke dataset.
The rest of the paper is organized as follows: Section II provides a brief overview of the related literature. Section III introduces the fundamental concepts of $t\bar{a}la$ in Hindustani music. Section IV provides a detailed description of the proposed method for TST, followed by $t\bar{a}la$ identification, and finally, ADT. Section V outlines the experimental setup for evaluating the proposed approach, including the datasets used for TST, $t\bar{a}la$ identification, and ADT. Section VI discusses the experimental evaluation and findings, and Section VII concludes the paper with key insights and future directions.
II Literature review
II-A TST and ADT
Significant work has been done on ADT in Western music [1]. Recent ADT methods utilize deep learning with various model architectures and feature representations [10, 11, 12, 1, 13, 14, 15, 16, 17, 18, 19, 2]. ADT identifies stroke onsets, marked by sudden energy increases. The attack-decay-sustain-release (ADSR) sequence forms the general temporal structure of a stroke. Detecting strokes in a drum kit is relatively straightforward due to distinct timbral structures and variations in shape and material [1]. Drum strokes feature sharp attacks, fast decay, and uniform repetition [20]. In contrast, TST is more challenging than ADT because of the significant timbral similarities and irregular shapes of strokes produced by the same instruments. This complexity requires sophisticated techniques for accurate transcription, making rhythmic analysis in Hindustani music more intricate. This work focuses on TST, emphasizing $t\bar{a}la$ identification and further validation of TST methods alongside ADT.
The literature on TST employs two main approaches: segment-classify and deep learning. The segment-classify method detects stroke onsets using thresholding on spectral flux, then classifies segments with various classifiers. Studies [21, 22, 23, 24, 25, 26, 27] follow this approach. Early work used Gaussian Mixture Models (GMM) and Hidden Markov Models (HMM), later extended with spectral and temporal features for neural networks and SVM classifiers [21]-[23]. Studies [24]-[27] focused on acoustic features and ADSR properties with MLP and SVMs. The DL approach involves spectrograms or Mel spectrograms processed through networks for transcription, with recent studies [9, 5, 6] employing CNNs, data augmentation, and transfer learning techniques.
Early studies [21, 22, 23, 24, 25, 26, 27] trained models on limited datasets, often from a single tabla. Recent studies [5, 6] used more realistic datasets with harmonium accompaniment but still employed only four stroke classes, insufficient for actual $t\bar{a}la$ identification. Additionally, these models have not been tested on concert data with vocals and other instruments, raising doubts about their performance on actual stroke classes and concert datasets.
II-B Meta-learning
Meta-learning has been extensively explored in image processing and computer vision to enable deep-learning models with minimal data samples. However, its application in the audio domain remains comparatively limited. Meta-learning mimics the human ability to learn from previous experiences or knowledge. Meta-learning algorithms are further divided into metric-based, model-based, and optimization-based approaches. References [28, 29, 30, 31, 32, 33, 34, 35, 36] primarily focus on metric-based few-shot learning, especially using prototypical networks. For instance, [30, 36] applies prototypical networks to drum transcription, treating each instrument as a binary classification task. Similarly, [31, 32, 33] handle individual audio event clips as binary classification tasks. In [34], multi-class audio segments are framed as binary classification, with the target class as positive and others as negative. Furthermore, [35] adapts the MAML algorithm for sound localization in new room environments, while [37] applies it to melody extraction.
II-C $T\bar{a}la$ identification (Music sequence matching)
String-matching techniques have been extensively explored in the literature for comparing two musical fragments. In [8], Longest Common Subsequence (LCS) and Rough Longest Common Subsequence (RLCS) matching methods are presented. In [38], a modified version of RLCS is discussed for detecting motifs within $alapanas$ in Carnatic music. Similarly, [9] and [39] employ a modified RLCS approach for searching specific stroke patterns in tabla stroke transcription. In [40], $R\bar{a}ga$ verification utilises the Longest Common Segment Set (LCSS). However, LCS, RLCS, and LCSS operate on the subsequence principle, where elements occur in the same order but not necessarily contiguously. In the context of $t\bar{a}las$, where continuity and order are crucial, these measures may not be suitable for analysing continuous sequences. References [3] and [41] focus on matching beat amplitude patterns, where a specific stroke corresponds to a particular amplitude. However, this approach may fail in practical situations where beat amplitudes do not consistently follow the same pattern.
Figure 1: System overview for automatic $t\bar{a}la$ identification
III $T\bar{a}la$ in Hindustani Music
In Hindustani music, a $t\bar{a}la$ describes the rhythmic framework of any composition. A $t\bar{a}la$ consists of a fixed time cycle divided into equal primary time units known as $m\bar{a}tra$. These $m\bar{a}tra$ are grouped into sections, sometimes with varying durations, referred to as $vibh\bar{a}gs$. The complete cycle is called an $avart$, and the start of the cycle (the downbeat) is called $s\bar{a}m$. Each $t\bar{a}la$ features pre-defined rhythmic patterns called the $thek\bar{a}$ [42]. Comprehensive explorations of $t\bar{a}la$ in Hindustani music can be found in [43, 42]. The tabla is the main percussion instrument that acts as the timekeeper of $t\bar{a}la$. It comprises two drums: the left-hand bass drum, known as the $bayan$ or $dagga$, and the right-hand drum, referred to as the $dayan$ or $tabla$. These drums can produce various pitched sounds and are played using palms and fingers.
[…]
Even though Hindustani and Western music differ, comparing some terms is useful. In Hindustani music, $m\bar{a}tra$ and $avart$ are similar to ‘beat’ and ‘bar’ in Western music. To clarify things, we’ll use ‘beat’ instead of ‘$m\bar{a}tra$’ and ‘stroke’ instead of ‘$b\bar{o}l$.’ This helps us understand the rhythms better across different musical styles.
IV Proposed Method
This article introduces a MAML-based approach for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT) with limited data. By leveraging the stroke sequence obtained from TST, we perform $t\bar{a}la$ identification using two proposed methods. The proposed tala identification system is illustrated in Fig. 1, showcasing the integration of these components.
IV-A Tabla Stroke Transcription (TST)
This section presents our solution for tabla stroke transcription with minimal labeled data. We train and test deep learning models using three methods to identify the optimal transcription approach. Subsequently, we leverage the pre-trained model weights from the best method for transfer learning and model-agnostic meta-learning (MAML) for tabla stroke transcription.
IV-A1 Step1: Pre-training
In this context, we implement three distinct transcription methods, as explained below.
One-way Transcription Method (OTM) [5, 6]
We adopt the method introduced by Rohit et al. [5, 6], expanding the stroke classes from four to ten. Ten distinct CNN models are employed to classify the ten stroke classes. Let $(X_{i},Y_{i})$ represent a single sample pair from a labeled dataset comprising $N$ samples of individual strokes. Here, the input $X_{i}\in\mathbb{R}^{3\times F\times T}$ to the model consists of a 3-channel chunk of 15-time frames $(T)$ extracted from a stack of 3 log-Mel spectrograms. The model $f_{\theta}$, parameterized by $\theta$, predicts the presence of onset for the center frame by leveraging contextual information from surrounding frames. It compares its output against a binary target $Y_{i}\in{0,1}$, where 1 denotes stroke onset, and 0 denotes no onset.
Proposed Transcription Method - 1 (PTM1)
This work proposes a new mridangam and tabla stroke transcription method inspired by Sound Event Detection (SED) tasks [44]. The proposed method utilizes a CRNN for stroke transcription. The labeled dataset of $N$ samples is denoted as $D(X_{i},Y_{i})$, where $i=0,1,…,N$. Here, $X_{i}\in\mathbb{R}^{F\times T}$ is the input feature matrix (standardized log Mel-spectrogram), containing $F$ acoustic features extracted from each frame in each audio recording with $T$ time frames. The labels are one-hot encoded as the target output matrix $Y_{i}\in\left\{0,1\right\}^{C\times T}$, where $C$ represents the number of predefined stroke classes.
Let $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ be the multi-class classification model that maps the given input data $X$ to the target outputs $Y$. Here, $\theta_{1}$, $\theta_{2}$, and $\theta_{3}$ are the parameters of convolutional, recurrent, and classifier layers, respectively. For a given feature vector $x$ for a single time frame, $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ outputs the stroke label with the highest soft-max probability. During the training phase, the model parameters $\theta_{1}$,$\theta_{2}$, and $\theta_{3}$ are randomly initialized and updated using the stochastic gradient descent algorithm as:
$[\theta_{1},\theta_{2},\theta_{3}]\leftarrow[\theta_{1},\theta_{2},\theta_{3}]-\alpha\nabla_{[\theta_{1},\theta_{2},\theta_{3}]}L_{wCE}\left(f_{[\theta_{1},\theta_{2},\theta_{3}]}\right)$ (1)
where, $\alpha\in\mathbb{R}^{+}$ is the learning rate, and $L_{wCE}$ is weighted categorical cross-entropy loss to handle the class imbalance defined as:
$L_{wCE}=-\sum_{c=1}^{C}w_{c}y_{i}\ log(\hat{y_{i}})$ (2)
where, $C$ is the number of classes, $\hat{y_{i}}$ is the predicted output, and $y_{i}$ is the ground truth for $i^{th}$ time frame for class $c$, i.e., 0 or 1 and $w_{c}\in\mathbb{R}^{+}$ is inversely proportional to the total number of time frames in training data corresponding to each class $c$.
Post-processing is applied to these predictions in two steps: First, if the predicted class label of $frame_{i-1}$ and $frame_{i+1}$ are the same but different from the class label of frame $frame_{i}$ then the class label of $frame_{i}$ is changed to match $frame_{i-1}$ and $frame_{i+1}$. In the second step, the frame where the class label changes from one class to another is identified as the onset frame for the new class. All onset frames in the test sample are identified, and the temporal locations of the onsets are calculated from these frames.
Proposed Transcription Method - 2 (PTM2)
[…]
IV-A2 Step 2 - A): Model Agnostic Meta-learning (MAML)
The complete MAML process can be divided into meta-training and meta-testing phases. The model is trained on various tasks during meta-training to adapt to new tasks effectively with only a few samples. Additionally, the model is evaluated on different distribution/domain datasets along with different or same tasks in the meta-testing phase to assess its adaptability to unseen, diverse scenarios.
Meta-training
Let, a task $T$ is sampled from a distribution $p(T)$ within the meta-training dataset $D_{train}$. Each task $T_{i}\sim p(T)$ is divided into a support set $T_{i}^{s}$ consisting of $s$ samples and a query set $T_{i}^{q}$ containing $q$ samples. The pre-trained base model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$, with parameters $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ explained in Section IV-A1, serves as the base-learner model. During meta-learning, we freeze the parameters of the convolutional layers and update the parameters of the recurrent and classifier layers. Thus, $\theta_{1}$ remains the same as in the pre-trained model while $\theta_{2}$ and $\theta_{3}$ become trainable parameters. We denote the combined parameters $\theta_{2}$ and $\theta_{3}$ as $\phi$, making $\phi$ the meta parameters of the base-learner model. The updated parameters of the base learner after $N$ steps for task $T_{i}$ are given by,
$\phi_{N}^{i}=\phi_{N-1}^{i}-\alpha\nabla_{\phi}L_{T_{i}^{s}}\left(f_{[\theta_{1},\phi_{N-1}^{i}]}\right)$ (3)
where, $\alpha$ represents the base model’s learning rate, and $L_{T_{i}^{s}}(f_{[\theta_{1},\phi_{N-1}^{i}]})$ denotes the loss computed on the support set of task $T_{i}$ after $(N-1)$ update steps, as given in Equation (2). This process, known as inner loop optimization (ILO), involves updating the model $f_{[\theta_{1},\phi]}$ based on the support set. Once base learning is complete, the model $f_{[\theta_{1},\phi]}$ becomes $f_{[\theta_{1},\phi_{N}^{i}]}$. Subsequently, using these updated parameters $f_{[\theta_{1},\phi_{N}^{i}]}$, the query loss $L_{T_{i}^{q}}({\phi}_{N}^{i},T_{i}^{q})$ is computed on the query set $T_{i}^{q}$. The meta-parameters $\phi$ are then updated using the query loss. The process of updating meta-parameters over the batch of tasks is called outer loop optimization (OLO) and is given by,
$\phi\leftarrow\phi-\beta\nabla_{\phi}\sum_{T_{i}}L_{T_{i}^{q}}\left(f_{[\theta_{1},{\phi}_{N}^{i}]}\right)$ (4)
where $\beta$ is a meta-learning rate and $L_{T_{i}^{q}}$ is loss on the query set $T_{i}^{q}$ for task $T_{i}$ calculated by equation (2). The entire ILO and OLO process (two-stage optimization) is repeated for all tasks $T_{i}$ in the $D_{train}$ dataset for $E$ epochs. The complete training process of MAML is outlined in Algorithm 1.
Algorithm 1 MAML for stroke detection
0:  Pre-trained base model parameters $[\theta_{1},\phi]$; frozen $\theta_{1}$
0:  $\alpha$, $\beta$: learning rates
1:  for E number of epochs do
2:    for all tasks $i$ in dataset $D_{train}$ do
3:       Initialize $\phi^{i}=\phi$
4:       Sample a batch of Log Mel spectrograms as $T_{i}\sim p(T)$
5:       Select $s$ samples as a support set $T_{i}^{s}$ and      $q$ samples as a query set $T_{i}^{q}$
6:       Update base-learner parameters $\phi_{N}^{i}$ using support set      $T_{i}^{s}$ by ILO ($N$ update steps) given by equation (3)
7:       Update $\phi$ using query set $T_{i}^{q}$ by OLO (1 update step)     given by equation (4)
8:     end for
9:  end for
10:  Obtained updated parameters $\phi$
Meta-testing
We test the trained model $f_{[\theta_{1},\phi]}$ in this stage. The updated model parameters $\phi$ from the meta-training phase now serve as good initialization parameters for adapting to new stroke classes with few samples. Given a new task $T_{j}$ with new unseen stroke classes from the test data $D_{test}$ consisting of a support set $T_{j}^{s}$ with $s$ samples and a query set $T_{j}^{q}$ with $q$ samples. The model $f_{[\theta_{1},\phi^{j}]}$ is initialized with $\phi^{j}=\phi$ and trained on support set $T_{j}^{s}$ using equation (3). After $N$ update steps, the updated parameters become $\phi_{N}^{j}$. This ILO process is repeated for $E_{1}$ iterations. The model’s performance with final updated parameters $\phi_{N}^{j}$ is then evaluated on the query set $T_{j}^{q}$.
IV-A3 Step 2 - B): Transfer learning
We use pre-trained models from the mridangam dataset and fine-tune them with a small tabla stroke dataset to leverage their acoustic similarities. Transfer learning is applied to a pre-trained model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ with parameters $\theta_{1}$, $\theta_{2}$ and $\theta_{3}$, as described in Section IV-A1. During fine-tuning, $\theta_{1}$ remains fixed, while $\theta_{2}$ and $\theta_{3}$ are combined and updated as $\phi$. The parameters $\phi$ are optimized using the stochastic gradient descent algorithm Equation (1) and the weighted cross-entropy loss Equation (2).
IV-B $T\bar{a}la$ identification
$T\bar{a}las$ are primarily identified by their $th\bar{e}k\bar{a}$, which maintains a fixed stroke count as discussed in Section III. We utilize these sequence and ratio properties for $t\bar{a}la$ identification, defining two $t\bar{a}la$ identification methods (scores): ‘NW Matching Score’ and ‘Stroke Ratio Score,’ based on the Needleman-Wunsch algorithm [45] and Cosine Similarity, respectively.
Algorithm 2 Algorithm for NW Matching Score
1:  Let $X_{ref}=\left\langle x_{0},x_{1},...,x_{m-1}\right\rangle$ be a reference sequence and $Y=\left\langle y_{0},y_{1},...,y_{m-1}\right\rangle$ be a test sequence frame of $m$ strokes. $S\in\mathbb{Z}^{(m+1)X(m+1)}$ be NW Matching Score matrix, where $S[i][j]$ is the matching score of the first $i$ strokes of $X_{ref}$ with the first $j$ strokes of $Y$.
2:  Initialization:$S[0][0]=0$, $S[i][0]=S[i-1][0]+gap\ penalty$ , $\forall i=1,...,m$, $S[0][j]=S[0][j-1]+gap\ penalty$ , $\forall j=1,...,m$, where, $gap\ penalty$ is ‘$-2$’.
3:  NW Matching Score matrix, $\forall\ i,j=1,...,m$,
$S[i][j]=max\begin{cases}&\text{$S[i-1][j-1]+match\ score$}\\ &\text{$S[i-1][j]+gap\ penalty$}\\ &\text{$S[i][j-1]+gap\ penalty$}\\ \end{cases}$
where, the $match\ score$ is ‘$1$’ if stroke $x_{i-1}$ and $y_{j-1}$ are same and ‘$-1$’ if both are different
4:  The optimum score is the sum of the scores along the optimal path obtained by backtracking from $S[m][m]$ to $S[0][0]$.
Figure 2: The NW Matching Score compares a transcripted sequence of $n$ strokes with a reference sequence of $m$ strokes.
IV-B1 NW Matching Score
We obtain a stroke sequence $Y=\left\langle y_{0},y_{1},...,y_{n-1}\right\rangle$ of $n$ strokes from the tabla stroke transcription module. We compare it with a reference $t\bar{a}la$ sequence $X_{ref}=\left\langle x_{0},x_{1},...,x_{m-1}\right\rangle$ of $m$ strokes. Since the test audio can be taken from any part of the complete audio, it doesn’t need to start from the first stroke of $X_{ref}$ (i.e., $s\bar{a}m$). We use a frame-shifting approach, sliding an $m$-stroke frame over $n$ strokes, as shown in Fig. 2. The score calculation for the frame of $m$ stroke sequence is shown in Algorithm 2. We compute the maximum score of every $m$ shift. The final $t\bar{a}la$ matching score $\sigma_{nw}$ averages all maximum scores over $n$ strokes, given by
$\sigma_{nw}=\frac{\sigma_{nw_{0}}+\sigma_{nw_{1}}+...+\sigma_{nw_{k}}}{k}$ (5)
where $k$ are steps to cover transcription sequence. The computational details of $\sigma_{nw_{0}},\sigma_{nw_{1}},...,\sigma_{nw_{k}}$ are shown in Fig. 2. This approach addresses potential missing strokes in specific frames of the test sample.
IV-B2 Stroke Ratio Score
We obtain the stroke count from the tabla stroke transcription module and calculate the stroke ratio for the given audio file. We analyze the beat proportions in four $t\bar{a}las$ as follows: $T\bar{i}nt\bar{a}la$ with $[Dha,Dhin,Tin,Ta]=[3,3,1,1]$, $\bar{E}kt\bar{a}la$ with $[Dhin,Tun,Na,Kat,Ta,Dhage,Tirkita]=[3,1,2,1,1,2,2]$, $Jhapt\bar{a}la$ with $[Dhi,Na,Ti]=[5,4,1]$, and $R\bar{u}pak\ t\bar{a}la$ with $[Tin,Na,Dhi]=[2,3,2]$. To identify the best match, we compute the cosine similarity between the test audio stroke ratio and the reference ratios for each $t\bar{a}la$. The Stroke Ratio Score is calculated as follows,
$Stroke\ Ratio\ Score=\frac{R.T}{\left\|R\right\|\left\|T\right\|}$ (6)
where, $R$ is the reference stroke ratio vector, and $T$ is the test stroke ratio vector.
IV-C Automatic Drum Transcription (ADT)
This section presents our solution for ADT with minimal labelled data. We employ transfer learning and model-agnostic meta-learning (MAML) for ADT.
IV-C1 MAML for ADT
A similar MAML process, as explained in Section IV-A2, is adopted for ADT. This approach employs a CRNN model as the base learner for multi-label classification. The input feature matrix is a standardised log Mel-spectrogram, similar to those used in PTM1 and PTM2, with the target vector methodology following the approach described in [12, 13]. Unlike tabla stroke transcription, where the convolutional layer parameters are frozen during meta-training and meta-testing, all parameters $\theta_{1},\theta_{2},\theta_{3}$ are trainable in ADT. These parameters are optimised using the weighted binary cross-entropy loss and updated via the stochastic gradient descent algorithm.
IV-C2 Transfer Learning for ADT
A similar transfer learning approach, as detailed in Section IV-A3, is utilized for ADT. The model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ is initially trained on a source domain dataset and subsequently fine-tuned on a target domain dataset. Model parameters are optimized using the stochastic gradient descent algorithm with the weighted binary cross-entropy loss.
V Experimental Setup
This study employs simple deep CNN and CRNN models for mridangam and tabla stroke transcription and the ADT task across all experiments. The architectures of the CNN and CRNN models are depicted in Fig. 3. Comprehensive details regarding the datasets utilized and the experimental settings for each method are outlined in the following sections.
Figure 3: General CNN [6] and CRNN model architecture for mridangam and tabla stroke transcription
V-A Datasets
For ADT, we use the $D_{D1}$ : IDMT-SMT-Drums [46] (DTD type), $D_{D2}$: ENST-Drums minus-one [47] (DTP type), $D_{D3}$: MDB-Drums [48] (DTD and DTM type), and $D_{D4}$ : ADTOF [49] (DTM type).
For TST, we use the $D_{T1}$ : Tabla Solo Dataset [9], $D_{T2}$ : the 4way-tabla-ismir21-Dataset [5], and $D_{T3}$: the Hindustani Music Rhythm Dataset [43]. $D_{T1}$ and $D_{T2}$ are tabla-only datasets, while $D_{T3}$ includes tabla along with melodic instruments and vocals, forming a complete concert dataset. Additionally, we curated a Synthetic Mridangam Stroke Dataset $(D_{M})$, as explained below.
Synthetic Mridangam Stroke Dataset $(D_{M})$
We curated a dataset of 720 audio files, each with a duration of 10 seconds, by concatenating randomly selected mridangam strokes from the existing Mridangam Stroke Dataset [50]. This original dataset includes audio samples for ten distinct stroke labels across six different tonic pitches, recorded at a sampling rate of 44.1 kHz. The resulting dataset consists of 120 audio files for each tonic, amounting to a total audio duration of 120 minutes.
TABLE I: TST : Meta-train-test dataset
Train/Test Datasets Stroke Classes
Meta-train
$D_{T1}$
Da, Ki, Ge, Kda, Tit,
Dha, Dhe, Dhi, Re, Tin
Meta-test-1
$D_{T1}$ Ta, Na, Din, Dhet, Dhin
Meta-test-2
$D_{T2}$
B, D, RB, RT
Meta-test-3
$D_{T3}$
Dha, Dhin, Tin, Na, Tun, Kat,
Ta, Dhage, Tirkita, Dhi, Ti
TABLE II: ADT : Meta-train-test dataset
Train/Test Dataset Stroke classes
Meta-train $D_{D4}$ BD, SD, TT, HH, CY+RD
Meta-test-1 $D_{D1}$ KD, SD, HH
Meta-test-2 $D_{D2}$
BD, SD, CL, HH, BE,
TT, RD, CY
Meta-test-3 $D_{D3}$
KD, SD, CHH, HH, TT,
RD, CY, OT, TMB
V-B Tabla Stroke Transcription (TST)
V-B1 Step - 1: Pre-training on $D_{M}$
One-way Transcription Method
We employ 10 CNNs for 10 classes as discussed in V-A. Akshay et al. [51] categorized these classes into damped (D), resonant-treble (RT), resonant-bass (RB), and resonant-both (B). We use the same model architectures as Rohit et al. [5] for these ten classes as follows: for $Thi$, $Ta$, $Num$, $Tha-$D; for $Dhin$, $Cha$, $Bheem-$ RT; for $Thom-$RB; and for $Tham$, $Dheem-$B. To achieve tonic-independent inference, we applied 6-fold cross-validation, splitting the dataset into six folds by tonic. We trained models on five folds and tested on one, with each fold’s training data split 90:10 for training and validation. This process covered all six tonics and we averaged the evaluation scores.
The CNN processes log-Mel spectrograms from three channels with window durations of 23 ms, 46 ms, and 92 ms, all using a fixed hop size of 10 ms. Each channel’s input has dimensions of 128 Mel-frequency bands by 15 time frames. Training employs the Adam optimizer with a batch size 256 for up to 150 epochs. Early stopping with a patience of 10 epochs is used, saving the model with the lowest validation loss. During inference, a fixed-threshold peak picking algorithm [52] binarizes the output, identifying significant peaks to convert continuous values into binary decisions.
Proposed Transcription Method - 1 (PTM1)
The datasets used in this work vary in audio length, so we segment them into nonoverlapping 5-second chunks for feature extraction, each labeled accordingly. The input to the CRNN is a standardized log Mel spectrogram, computed with a 46.4 ms window size and a 10 ms hop size, utilizing 128 Mel filter banks. Each frequency band of the Mel spectrogram is normalized to have zero mean and unit variance. This method assigns equal importance to the duration of the entire stroke rather than focusing solely on the onset frame.
Mridangam and tabla strokes are primarily characterized by the attack, decay, and sustain phases of the ADSR model, with most strokes having a fast-decaying exponential envelope [21]. Therefore, stroke onset is predominantly defined by the energetic attack-decay phase, while the sustain phase defines how the stroke ends. The last part of the release region contributes minimally to defining stroke onset and class. If an equal focus is given to the complete stroke duration during transcription, it can confuse the model since this phase is similar across all stroke classes. Hence, we transcribe this segment separately as ‘No-stroke’ using a threshold set at 3% of the stroke’s maximum amplitude. Frames from onset to threshold denote the stroke class, while frames from threshold to the subsequent stroke onset are labeled ‘No-stroke.’ We employ 6-fold cross-validation using six tonics. Training involves a batch size of 32 over 100 epochs, with a learning rate of 0.001. During inference, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class for evaluation.
Proposed Transcription Method - 2 (PTM2)
The CRNN model architecture and normalized log Mel spectrogram input to the CRNN are the same as explained in Section V-B1. The training targets include the ‘No-stroke’ class, as detailed in Section IV-A1. The model is trained with a batch size of 32 for 100 epochs, using a learning rate of 0.001. During inference, the model predicts stroke classes frame-by-frame for a given test log-Mel spectrogram. The onset of each stroke class corresponds to the frame in which the model predicts it. Then, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class.
V-B2 Step 2
Based on the pretraining results, one-way transcription performs poorly in complexity and time consumption, requiring a CNN for each stroke class. As a result, we will discontinue this approach and instead focus on the proposed methods. Specifically, we will employ the CRNN model for MAML-based domain adaptation and transfer learning-based domain adaptation for tabla strokes. Our methodology will be validated across three tabla datasets.
MAML
We meta-train our models on ten stroke classes from $D_{T1}$ and conduct meta-testing in three scenarios: evaluating performance on five stroke classes from $D_{T1}$, testing with four stroke classes from $D_{T2}$, and assessing eleven stroke classes from $D_{T3}$. Additionally, we performed experiments using five, six, and eleven stroke classes from $D_{T3}$ to study the impact of class number in meta-testing. Detailed stroke class distributions for meta-training and meta-testing are in Table I. We randomly selected 12 minutes of audio from $D_{T1}$ for meta-training and used 5 minutes from $D_{T1}$ , 10 minutes from $D_{T2}$ , and 10 minutes from $D_{T3}$ for meta-testing. Furthermore, the model adapted during meta-testing on $D_{T2}$ was directly evaluated on a separate test partition from $D_{T2}$ provided by Rohit et al. [5], enabling comparison with the baseline approach.
In the first scenario, we must partition the existing classes into two mutually exclusive sets. Due to the challenge of directly segmenting long audio based on stroke classes, we adopt an approach from Nolasco et al. [28] and Shi Bowen et al. [34]. Here, we preserve the original class label for desired strokes and classify non-target strokes as ’Others.’ As a result, there will be differences in the classes used between meta-training and meta-testing.
Although some stroke types in $D_{T2}$ and $D_{T3}$ may appear similar, they are performed with different background instruments: $D_{T3}$ includes vocals and various instruments, while $D_{T2}$ is standalone. Therefore, we treat them as distinct stroke types. The ‘No-stroke’ class is also considered in the ‘PTM1’ and ‘PTM2’, discussed earlier in Sections IV-A1 and IV-A1, respectively. The MAML approach is applied to both methods with the same experimental settings, utilizing their respective input-output pairs. For task $T_{i}$, we randomly sample log Mel spectrograms with a batch size comprising 32 support samples and 8 query samples from the meta-training dataset. Models are trained with the Adam optimizer for 300 epochs, employing an inner loop optimization of $N=3$. The learning rates $\alpha$ and $\beta$ are fixed at 0.001.
Transfer Learning
We validate the transfer learning method using the same classes and samples from $D_{T1}$, $D_{T2}$, and $D_{T3}$ as in Section V-B2. Additionally, a fine-tuned model on $D_{T2}$ is tested on separate test data provided by Rohit et al. [5] in $D_{T2}$. Transfer learning is applied with the same experimental settings to both methods described in Sections IV-A1 and IV-A1, using their respective input-output pairs. The models are trained with a batch size of 32 for 100 epochs and a learning rate of 0.001.
V-C Automatic Drum Transcription (ADT)
V-C1 MAML for ADT
Meta-training is performed using five stroke classes from $D_{D4}$ while meta-testing is carried out in three distinct scenarios: (i) three stroke classes from $D_{D1}$, (ii) eight stroke classes from $D_{D2}$, and (iii) nine stroke classes from $D_{D3}$. The detailed distribution of stroke classes used for meta-training and meta-testing is provided in Table II. The complete $D_{D4}$ dataset is utilized for meta-training, and 5-minute excerpts from $D_{D1}$, $D_{D2}$, and $D_{D3}$ are used for meta-testing.
For evaluation, the model trained during meta-testing is directly tested on the remaining portions of each dataset to compare its performance against the baseline approach. The MAML framework follows the same experimental setup as in tabla stroke transcription. The state-of-the-art CRNN model from [15] is also trained and evaluated on the same meta-testing data to ensure a consistent and fair performance comparison.
V-C2 Transfer learning for ADT
In the first step, the CRNN model is trained on $D_{D4}$ and subsequently fine-tuned using the same classes and samples from the meta-test datasets $D_{D1}$, $D_{D2}$, and $D_{D3}$. The fine-tuned model is then evaluated on the remaining portions of each dataset, as detailed in Section V-C1. The training employs a batch size of 32 over 100 epochs, with a learning rate 0.001. This experimental setup thoroughly assesses the model’s transfer learning capabilities across different datasets.
VI Results and Discussions
TABLE III: Step-1: Transcription Performance on $D_{M}$ (f1-score)
Approach $Bheem$ $Cha$ $Dheem$ $Dhin$ $Num$ $Ta$ $Tha$ $Tham$ $Thi$ $Thom$ Average
OTM 83.2 47.3 55.6 51.6 76.4 64.8 70 73.3 61.1 67.1 65.04
PTM2 95.4 86.7 87.3 78.2 91.6 85.9 91.4 83.3 85.3 86.5 87.16
PTM1 96.6 93.6 94.1 92.5 94.6 92.2 95.1 89.9 93.4 89.7 93.17
  • OTM: One-way Transcription Method, PTM1:Proposed Transcription Method-1, PTM2:Proposed Transcription Method-2
Performance is evaluated using the f1-score with a 50 ms collar for detecting onset positions. Scores are computed for each stroke type on individual tracks and then averaged across the dataset using the $mir\_eval$ Python library for both TST and ADT.
VI-A Tabla stroke transcription (TST)
Table III shows the average cross-validation results of pre-training on $D_{M}$ using the three methods from Section IV-A1. Predictions are made by identifying and thresholding local peaks in the network’s output activations. The threshold is chosen to maximize the f-score on the validation set and then applied to the test set.
TABLE IV: Step-2: TST Performance on $D_{T1}$ (f1-score)
Approach $Other$ $Ta$ $Na$ $Din$ $Dhet$ $Dhin$ Average
PTM2+TL 69.1 48.7 50.4 51.9 58.1 53.1 55.22
PTM1+TL 80.1 58.9 54.9 57.7 59.9 61.5 62.17
PTM2+ML 82.4 58.8 69.8 89.3 91.4 67.5 76.53
PTM1+ML 87 71.2 76.1 89.9 92.1 71.6 81.32
TABLE V: Step-2: TST Performance on $D_{T2}$ (f1-score)
Approach B D RB RT Average
Drum-pretrained [6] 2.1 44 17.5 48.7 28.05
Best set of D, RT, RB [6]
81.5 83 63.6 86 78.52
Retrained [6] 82.7 83.6 66.9 86.6 79.95
Retrained Rohit [5]
80.1 83.3 34.1 84.3 70.4
PTM2 + TL 57.9 59 35.6 65.6 54.52
PTM1 + TL 59.4 72.1 38.6 71 60.28
PTM2 + ML 81 79.6 63.7 82.9 76.8
PTM1 + ML 87.3 85.8 69.4 89.9 83.1
TABLE VI: Step-2: TST Performance on $D_{T3}$ (11 stroke classes) (f1-score)
Approach $Dha$ $Dhin$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Average
PTM2+TL 19.2 36.9 9.9 34.8 16.3 21.3 26.4 39.2 27.9 22.8 23.1 25.25
PTM1+TL 24.3 45.8 24 50.7 61.5 26.8 48.7 47.6 45.2 33.5 26.1 39.47
PTM2+ML 39.4 61.2 31.3 45.5 49 65.2 65.6 64.1 61.2 46.9 38.1 51.59
PTM1+ML 45.5 66.9 38 55.2 87.8 71.6 81.4 78.2 78.7 49.5 40.5 63.02
  • PTM1:Proposed Transcription Method-1, PTM2:Proposed Transcription Method-2, TL:Transfer Learning, ML: Meta Learning
TABLE VII: Step-2: TST Performance on $D_{T3}$ (with 5 stroke classes) (f1-score)
Approach $Other1$ $Dhin$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Average
PTM2+TL 21.6 38.3 33.8 26.9 15.4 18.2 25.70
PTM1+TL 35.8 43.4 45.7 48.1 34.6 22.3 38.31
PTM2+ML 50.4 61.9 57.1 59.8 45.6 37.9 52.11
PTM1+ML 61.4 68.8 75.7 79.7 47.6 43.9 62.85
TABLE VIII: Step-2: TST Performance on $D_{T3}$ (with 6 stroke classes) (f1-score)
Approach $Other2$ $Dha$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ Average
PTM2+TL 31.5 17.2 13 37.3 19.4 24.6 28.9 24.55
PTM1+TL 42.3 19.5 22.6 47.7 59.1 25.8 50.3 38.19
PTM2+ML 54.1 40.1 29.7 47.7 48.9 66.3 67.4 50.6
PTM1+ML 60.1 46.2 36.7 56.8 85.9 72.3 82.4 62.91
TABLE IX: $T\bar{a}la$ identification performance on $D_{T3}$
Method Accuracy (%) Time (ms)
RLCS0[9] 33.6 64.2
NW Matching Score 48.9 62.77
Stroke Ratio Score 43.1 1.41
TABLE X: ADT performance on DTD - $D_{D1}$
(f1-score)
Approach KD SD HH Average
SOTA CRNN [15] 42.1 65.3 58.8 55.4
CRNN + TL 54.3 68.7 63.9 62.3
CRNN + ML 73.9 84.2 82.1 80.1
TABLE XI: ADT performance on DTP - $D_{D2}$ (f1-score)
Approach BD SD CL HH BE TT RD CY Average
SOTA
CRNN [15]
54.4 53.1 41.5 56.5 34.7 42.6 47.5 39.1 46.2
CRNN + TL 68.7 65.3 52.1 64.4 48.3 61.5 58.4 53.5 59.0
CRNN + ML 83.5 82.8 72.1 85.3 68.5 77.2 78 73.9 77.6
TABLE XII: ADT performance on DTD - $D_{D3}$ (f1-score)
Approach KD SD CHH HH TT RD CY OT TMB Average
SOTA CRNN [15] 44.8 46.3 38.7 33.1 27.4 39.3 32.6 24.5 14.4 33.4
CRNN + TL 54.3 70 63.5 60.2 57.5 62.7 54.1 44.8 33.4 55.6
CRNN + ML 73.1 91.6 84.4 82.3 70.9 81.6 74.5 67.3 59.5 76.1
TABLE XIII: ADT performance on DTM - $D_{D3}$ (f1-score)
Approach KD SD CHH HH TT RD CY OT TMB Average
SOTA CRNN [15] 41.7 38.2 29.8 30.3 24.5 36.2 29.4 18.6 8.6 28.6
CRNN + TL 52.7 66.1 61.7 57.8 53 60.3 50.1 40.9 31.8 52.7
CRNN + ML 71.4 87.8 80.1 79.8 68.6 78.5 64.9 61.3 54.2 71.8
TABLE XIV: ADT results Comparison
(Proposed vs. [36])
Approach % f1-score
Trained on SLAKH [36] 50.0
Trained on STAR [36] 62.0
Dynamic FSL [36] 70.0
CRNN + ML 71.8
The results demonstrate that the proposed methods outperform the existing OTM. The OTM, utilizing a CNN, only considers the spatial dimension of the Mel spectrogram. In contrast, the proposed CRNN method integrates both spatial and temporal dimensions, capturing the attack-decay-sustain-release (ADSR) temporal model of strokes more effectively. PTM1, which utilizes all stroke frames, outperforms PTM2, which relies on a single frame. PTM2’s reliance on a single frame may lead to missed stroke detections if the initial frame is undetected, whereas PTM1 remains robust by considering subsequent frames.
Table V shows meta-learning and transfer learning results using PTM1 and PTM2 on test data $D_{T1}$ from the same domain, demonstrating PTM1 with meta-learning outperforms PTM2 and others. Similarly, Table V presents results on test data $D_{T2}$ from a similar domain, with PTM1 again outperforming PTM2. Tables VI to VIII display results on vastly different domain test data $D_{T3}$, confirming PTM1 with meta-learning consistently outperforms PTM2 across different class variations.
Tables V-VIII demonstrate that the meta-learning consistently outperforms transfer learning. MAML learns an initialization for rapid task adaptation through meta-learning, while transfer learning directly uses pre-trained weights, which may not optimize performance for the target task. Specifically, Table V shows PTM1 with meta-learning outperforming four OTM settings using transfer learning and data augmentations, attributed to CNNs’ limited ability to capture temporal relationships. Tables VI-VIII indicate performance degradation on $D_{T3}$ due to frequent tabla stroke overlap with vocals and other instruments in concert data, complicating stroke onset detection. The number of stroke classes tested shows minimal impact on performance across these analyses.
VI-B $T\bar{a}la$ identification
After TST, $t\bar{a}la$ identification involves generating $t\bar{a}la$ identification scores. The average processing times for this task from transcribed tabla stroke sequences are 1.41 ms for cosine similarity scoring and 62.77 ms for sequence matching score when analyzing 2 minutes of test audio across 4 $t\bar{a}las$. These computations were performed on a system equipped with 16GB RAM, an Intel i7 processor, and a 6GB GPU.
VI-C Automatic drum transcription (ADT)
The performance of the proposed Transfer Learning and MAML-based approaches for ADT tasks is summarized in Tables X-XIII. These tables present results on datasets representing all ADT task types--DTD, DTP, and DTM--specifically $D_{DD1}$, $D_{DD2}$, and $D_{DD3}$. The comparison is made against the state-of-the-art CRNN model [15]. Additionally, Table XIV highlights the comparative performance of the proposed MAML approach with existing methods on $D_{D3}$ (DTM).
The results demonstrate that the proposed MAML-based method consistently outperformed existing approaches utilizing minimal target data across all ADT task types. This indicates the effectiveness of the proposed approach in scenarios with limited data availability.
VII Conclusion
This paper presents a novel approach leveraging Model-Agnostic Meta-Learning (MAML) for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT), addressing the challenge of limited annotated datasets. The proposed method effectively transcribes tabla strokes and drum sounds across diverse datasets, including scenarios with drum or tabla strokes alone, percussion instruments with melodic accompaniments, and the singer’s voice, ensuring comprehensive validation and demonstrating robustness and adaptability.
Following TST, two new $t\bar{a}la$ identification methods are introduced, contributing to rhythmic analysis in Hindustani music. Experimental results highlight the superiority of the MAML-based approach over existing methods, consistently achieving better performance with minimal labelled data.
Future work will focus on improving transcription accuracy for concert datasets with minimal labeled data, leveraging advanced music source separation techniques and active learning strategies.

### v3

@@
 This study introduces a meta-learning-based approach for low-resource Tabla Stroke Transcription (TST) and $t\bar{a}la$ identification in Hindustani classical music.
-Using Model-Agnostic Meta-Learning (MAML), we address the challenge of limited annotated datasets, enabling rapid adaptation to new tasks with minimal data.
+Using Model-Agnostic Meta-Learning (MAML), we address the challenges of limited annotated datasets and label heterogeneity, enabling rapid adaptation to new tasks with minimal data.
 The method is validated across various datasets, including tabla solo and concert recordings, demonstrating robustness in polyphonic audio scenarios.
@@
 AMT supports critical tasks such as melody extraction, chord recognition, beat tracking, and rhythm analysis, forming the foundation for advanced music research and applications [2].
-While substantial progress has been made in transcribing melodic instruments like the piano and guitar, the transcription of percussive instruments remains under-explored [1].
+While substantial progress has been made in transcribing melodic instruments such as the piano, guitar, strings, and winds, the transcription of percussive instruments remains relatively under-explored [1, 3].
 Furthermore, research progress in the percussion transcription of Hindustani classical music has been significantly slower compared to Western music [4].
@@
 It helps maintain the structure and essence of the chosen raga (melodic framework).
-Tabla Stroke Transcription (TST) is crucial for analysing the complex rhythmic patterns ($t\bar{a}las$) that form the basis of Hindustani music.
-Accurate TST also aids in preserving and digitising traditional music and creating learning tools for learners and musicians.
-Like Automatic Drum Transcription (ADT) in Western music, which focuses on drum sound transcription to understand rhythm structures, organise music collections and develop advanced tools, TST aims to identify and classify rhythmic events.
+Tabla Stroke Transcription (TST) is crucial for analyzing the complex rhythmic patterns ($t\bar{a}las$) that form the basis of Hindustani music.
+Accurate TST also aids in preserving and digitizing traditional music and creating learning tools for learners and musicians.
+Like Automatic Drum Transcription (ADT) in Western music, which focuses on drum sound transcription to understand rhythm structures, organize music collections and develop advanced tools, TST aims to identify and classify rhythmic events.
 However, unlike ADT, TST presents distinct challenges due to the tabla’s unique physical and timbral characteristics compared to drums [1, 5].
@@
 In Western music, bass guitars may overlap with bass drums, while melodic instruments or vocals can obscure tabla strokes in Hindustani music.
-Transcription methods address these challenges at different levels of complexity: Tabla/Drum-only Transcription (DTD) focuses on isolated recordings, avoiding polyphony.
-Tabla/Drum Transcription with Percussion (DTP) includes additional percussive elements, adding complexity.
-The most challenging, Tabla/Drum Transcription with Melodic Instruments (DTM), handles full audio mixes with all instruments and vocals.
+Transcription methods address these challenges at different levels of complexity: Drum/Tabla -only Transcription (DTD) focuses on isolated recordings, avoiding polyphony.
+Drum/Tabla Transcription with Percussion (DTP) includes additional percussive elements, adding complexity.
+The most challenging, Drum/Tabla Transcription with Melodic Instruments (DTM), handles full audio mixes with all instruments and vocals.
 Since DTM represents most real-world musical tracks, it is the most critical and complex task in transcription.
-Most of the present works in ADT rely on deep learning (DL) models due to their ability to handle complex audio mixtures.
-Similarly, few studies have been conducted on TST, with recent efforts also focusing on DL models [5, 6].
+Most existing ADT and recent TST methods rely on deep learning (DL) models due to their ability to handle complex audio mixtures [6, 7].
 However, DL models require large, annotated datasets for supervised training, which are scarce.
-Creating these datasets is challenging because annotations must be precisely aligned with the audio, a labour-intensive process requiring skilled musicians.
-Additionally, many annotated tracks are copyrighted, limiting their public availability.
-To address this, synthetic data generated from MIDI tracks has been used to produce large-scale training datasets with accurate labels.
-Despite their large volume, synthetic datasets often perform less effectively than real-world annotated data, making dataset scarcity a significant challenge for DL-based ADT and TST.
-To overcome this limitation, this paper introduces a new stroke transcription method based on model-agnostic meta-learning (MAML) [7], which operates effectively with minimal data.
-TST provides tabla strokes, and $t\bar{a}la$ is typically identified by stroke sequence.
-While some studies have explored music sequence matching [8, 9], these approaches are not directly applicable to $t\bar{a}la$ identification.
-In this study, we propose two new $t\bar{a}la$ identification methods.
-The key contributions of this work are summarised as follows:
+Creating these datasets is challenging because annotations must be precisely aligned with the audio, a labour-intensive process requiring skilled musicians, and many annotated tracks are copyrighted, limiting their public availability.
+Synthetic data generated from MIDI tracks has been used to produce large-scale training datasets with accurate labels, but such datasets often perform less effectively than real-world annotated data, making dataset scarcity a significant barrier for DL-based ADT and TST.
+Beyond low-resource conditions, a second challenge is label heterogeneity: stroke labels in tabla and drum datasets differ due to performance styles, labelling schemes, and musical traditions.
+For example, tabla datasets may follow $ghar\bar{a}n\bar{a}$ conventions, while Western drum datasets use distinct ontologies.
+These divergences lead to task heterogeneity with mismatched label spaces, reducing the reliability of standard transfer learning approaches that assume shared label ontologies and often requiring time-consuming label harmonization.
+Model-Agnostic Meta-Learning (MAML) enables fast adaptation across tasks with distinct label spaces [8, 9, 10, 11].
+This paper presents an MAML-based stroke transcription method addressing data scarcity and label inconsistency.
+In this work, we address the challenge of performing accurate Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT) with limited annotated data.
+Based on the transcribed stroke sequences from TST, we propose methods for $t\bar{a}la$ identification, as $t\bar{a}la$ forms the rhythmic backbone of Hindustani music.
+Identifying the $t\bar{a}la$ enables a deeper analysis of Hindustani music, and existing music sequence matching approaches [12, 13] are not directly applicable in this context.
+The key contributions of this work are summarized as follows:
   1. 1.
-    A novel meta-learning-based approach for Tabla Stroke Transcription (TST) is proposed and validated on both tabla solo and complete concert recordings.
-This method is further validated for Automatic Drum Transcription (ADT) tasks using DTD, DTP, and DTM datasets.
+    A novel meta-learning-based approach for TST is proposed and validated on both tabla solo and full concert recordings.
+This method is also extended to ADT tasks using the DTD, DTP, and DTM datasets, demonstrating its generalization across datasets.
   2. 2.
@@
     A synthetic dataset is curated from the existing mridangam stroke dataset.
-The rest of the paper is organized as follows: Section II provides a brief overview of the related literature.
-Section III introduces the fundamental concepts of $t\bar{a}la$ in Hindustani music.
-Section IV provides a detailed description of the proposed method for TST, followed by $t\bar{a}la$ identification, and finally, ADT.
-Section V outlines the experimental setup for evaluating the proposed approach, including the datasets used for TST, $t\bar{a}la$ identification, and ADT.
-Section VI discusses the experimental evaluation and findings, and Section VII concludes the paper with key insights and future directions.
+The rest of the paper is organized as follows: Section II provides an overview of related literature, Section III details the proposed method for TST and ADT, Section IV presents information about the datasets used, Section V outlines the experimental setup, Section VI discusses the experimental evaluation and findings, and Section VII concludes the paper with key insights and future directions.
+The dataset and supplementary materials are available at https://github.com/madhavlab/2025_drumtrans_rkodag
 II Literature review
@@
 Significant work has been done on ADT in Western music [1].
-Recent ADT methods utilize deep learning with various model architectures and feature representations [10, 11, 12, 1, 13, 14, 15, 16, 17, 18, 19, 2].
+Recent ADT methods utilize deep learning with various model architectures and feature representations [14, 15, 1, 16, 17, 18, 19, 2].
 ADT identifies stroke onsets, marked by sudden energy increases.
@@
 Detecting strokes in a drum kit is relatively straightforward due to distinct timbral structures and variations in shape and material [1].
-Drum strokes feature sharp attacks, fast decay, and uniform repetition [20].
-In contrast, TST is more challenging than ADT because of the significant timbral similarities and irregular shapes of strokes produced by the same instruments.
-This complexity requires sophisticated techniques for accurate transcription, making rhythmic analysis in Hindustani music more intricate.
+Drum strokes exhibit sharp attacks, rapid decay, and consistent repetition [20].
+In contrast, TST is more complex due to timbral similarities and variable stroke shapes from the same instrument, necessitating more sophisticated transcription methods and making rhythmic analysis in Hindustani music more intricate.
 This work focuses on TST, emphasizing $t\bar{a}la$ identification and further validation of TST methods alongside ADT.
@@
 The segment-classify method detects stroke onsets using thresholding on spectral flux, then classifies segments with various classifiers.
-Studies [21, 22, 23, 24, 25, 26, 27] follow this approach.
-Early work used Gaussian Mixture Models (GMM) and Hidden Markov Models (HMM), later extended with spectral and temporal features for neural networks and SVM classifiers [21]-[23].
-Studies [24]-[27] focused on acoustic features and ADSR properties with MLP and SVMs.
+Studies [21, 22, 23, 24] follow this approach.
+Early work used Gaussian Mixture Models (GMM) and Hidden Markov Models (HMM), later extended with spectral and temporal features for neural networks and SVM classifiers [21, 22, 23].
+Study [24] focuses on the acoustic features and ADSR properties of tabla strokes.
 The DL approach involves spectrograms or Mel spectrograms processed through networks for transcription, with recent studies [25, 6, 7] employing CNNs, data augmentation, and transfer learning techniques.
-Early studies [21, 22, 23, 24, 25, 26, 27] trained models on limited datasets, often from a single tabla.
+Early studies [21, 22, 23, 24] trained models on limited datasets, often from a single tabla.
 Recent studies [6, 7] used more realistic datasets with harmonium accompaniment but still employed only four stroke classes, insufficient for actual $t\bar{a}la$ identification.
@@
 II-B Meta-learning
-Meta-learning has been extensively explored in image processing and computer vision to enable deep-learning models with minimal data samples.
+Meta-learning, or “learning to learn,” is a training paradigm in which a model is optimized to rapidly adapt to new tasks using limited data [8].
+It mimics the human ability to transfer prior experience to new situations by learning how to learn from diverse tasks.
+Unlike traditional models that train on a fixed dataset and generalize within that domain, meta-learning involves exposure to a distribution of tasks during training, enabling the model to extract transferable knowledge.
+This knowledge allows the model to quickly fine-tune to new tasks with limited data.
+The approach typically involves two components: a base-learner that learns task-specific patterns, and a meta-learner that captures cross-task learning strategies to guide fast adaptation [8, 9, 10, 11].
+Meta-learning has been extensively explored in image processing and computer vision to enable deep-learning models with limited data samples.
 However, its application in the audio domain remains comparatively limited.
@@
 Meta-learning algorithms are further divided into metric-based, model-based, and optimization-based approaches.
-References [28, 29, 30, 31, 32, 33, 34, 35, 36] primarily focus on metric-based few-shot learning, especially using prototypical networks.
+References [26, 27, 28, 29, 30, 31, 32] primarily focus on metric-based few-shot learning, especially using prototypical networks.
 For instance, [27, 32] applies prototypical networks to drum transcription, treating each instrument as a binary classification task.
@@
 In [31], multi-class audio segments are framed as binary classification, with the target class as positive and others as negative.
-Furthermore, [35] adapts the MAML algorithm for sound localization in new room environments, while [37] applies it to melody extraction.
+Furthermore, [33] applies the MAML algorithm to melody extraction.
 II-C $T\bar{a}la$ Identification (Music Sequence Matching)
-String-matching techniques have been extensively explored in the literature for comparing two musical fragments.
-In [8], Longest Common Subsequence (LCS) and Rough Longest Common Subsequence (RLCS) matching methods are presented.
-In [38], a modified version of RLCS is discussed for detecting motifs within $alapanas$ in Carnatic music.
-Similarly, [9] and [39] employ a modified RLCS approach for searching specific stroke patterns in tabla stroke transcription.
-In [40], $R\bar{a}ga$ verification utilises the Longest Common Segment Set (LCSS).
-However, LCS, RLCS, and LCSS operate on the subsequence principle, where elements occur in the same order but not necessarily contiguously.
-In the context of $t\bar{a}las$, where continuity and order are crucial, these measures may not be suitable for analysing continuous sequences.
-References [3] and [41] focus on matching beat amplitude patterns, where a specific stroke corresponds to a particular amplitude.
-However, this approach may fail in practical situations where beat amplitudes do not consistently follow the same pattern.
-Figure 1: System overview for automatic $t\bar{a}la$ identification
-III $T\bar{a}la$ in Hindustani Music
-In Hindustani music, a $t\bar{a}la$ describes the rhythmic framework of any composition.
-A $t\bar{a}la$ consists of a fixed time cycle divided into equal primary time units known as $m\bar{a}tra$.
-These $m\bar{a}tra$ are grouped into sections, sometimes with varying durations, referred to as $vibh\bar{a}gs$.
-The complete cycle is called an $avart$, and the start of the cycle (the downbeat) is called $s\bar{a}m$.
-Each $t\bar{a}la$ features pre-defined rhythmic patterns called the $thek\bar{a}$ [42].
+In Hindustani music, a $t\bar{a}la$ defines the rhythmic structure of a composition.
+It comprises a fixed time cycle divided into equal time units called $m\bar{a}tras$, grouped into sections ($vibh\bar{a}gs$), forming a complete cycle known as an $avart$.
+The starting point of this cycle is referred to as $s\bar{a}m$, and each $t\bar{a}la$ features a characteristic rhythmic pattern known as the $thek\bar{a}$.
 Comprehensive explorations of $t\bar{a}la$ in Hindustani music can be found in [34, 35].
-The tabla is the main percussion instrument that acts as the timekeeper of $t\bar{a}la$.
-It comprises two drums: the left-hand bass drum, known as the $bayan$ or $dagga$, and the right-hand drum, referred to as the $dayan$ or $tabla$.
-These drums can produce various pitched sounds and are played using palms and fingers.
-[…]
-[…]
-Even though Hindustani and Western music differ, comparing some terms is useful.
-In Hindustani music, $m\bar{a}tra$ and $avart$ are similar to ‘beat’ and ‘bar’ in Western music.
-To clarify things, we’ll use ‘beat’ instead of ‘$m\bar{a}tra$’ and ‘stroke’ instead of ‘$b\bar{o}l$.’ This helps us understand the rhythms better across different musical styles.
-IV Proposed Method
-This article introduces a MAML-based approach for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT) with limited data.
-By leveraging the stroke sequence obtained from TST, we perform $t\bar{a}la$ identification using two proposed methods.
-The proposed tala identification system is illustrated in Fig.
-1, showcasing the integration of these components.
-IV-A Tabla Stroke Transcription (TST)
-This section presents our solution for tabla stroke transcription with minimal labeled data.
+Despite fundamental differences between Hindustani and Western music, the term equivalence aids clarity.
+Accordingly, we adopt ‘beat’ for $m\bar{a}tra$ and ‘stroke’ for $b\bar{o}l$ to facilitate cross-cultural rhythmic analysis.
+The tabla, consisting of two drums: the left-hand $bayan$ and the right-hand $dayan$, is the primary percussion instrument used to render the $thek\bar{a}$.
+A wide range of $t\bar{a}las$ exists in Hindustani music, with $T\bar{i}nt\bar{a}l$, $Ek\bar{a}l$, $Jhapt\bar{a}l$, and $Rupak\ t\bar{a}l$ being the most popular and the focus of this study.
+Six major $ghar\bar{a}n\bar{a}s$: Delhi, Ajrada, Lucknow, Banaras, Farrukhabad, and Punjab, define distinctive tabla styles.
+These styles vary in stroke articulation, dynamics, ornamentation, and improvisational phrasing, though the structure and number of strokes in the $thek\bar{a}$ typically remain consistent.
+Accurate $t\bar{a}la$ identification is essential for tasks such as rhythm analysis, segmentation, and transcription.
+String-matching techniques have been extensively explored in the literature for music sequence comparison.
+In [12], Longest Common Subsequence (LCS) and Rough LCS (RLCS) are used for symbolic matching.
+A modified RLCS approach is presented in [36] for motif detection in Carnatic $al\bar{a}panas$.
+In the context of tabla, [25]employ a modified RLCS approach for matching specific stroke sequences in transcribed performances.
+In [37], $R\bar{a}ga$ verification is performed using Longest Common Segment Set (LCSS).
+While these techniques preserve element order, they do not enforce continuity, limiting their effectiveness for $t\bar{a}la$ patterns where ordered and contiguous strokes are crucial.
+Alternatively, [4] propose beat amplitude pattern matching, assuming a correlation between strokes and amplitude.
+However, performance variability in dynamics often renders these methods unreliable in real-world scenarios.
+Consequently, there remains a need for matching strategies that retain both continuity and order in stroke sequences while accommodating expressive variations in timing and dynamics.
+Figure 1: System overview: (A) TST: Pre-trained on mridangam data using OTM, PTM1, PTM2, and adapted via MAML.
+(B) $T\bar{a}la$ identification: Identifies $t\bar{a}la$ using transcribed strokes from TST and NW Matching/Stroke Ratio scores.
+(C) ADT: Applies the same meta-learning strategies to Western drum audio.
+III Proposed Method
+This section describes our meta-learning-based system for low-resource Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT).
+Fig.1 illustrates the system overview, highlighting MAML’s role in enabling few-shot adaptability.
+The system consists of three components: A) A CRNN-based TST model trained with MAML that transcribes audio into stroke sequences.
+B) A downstream $t\bar{a}la$ identification module that infers rhythmic structure from these sequences.
+C) Validation of the framework on a Western drums dataset for the ADT task, demonstrating adaptability across diverse drum data.
+We introduce two novel scoring methods for $t\bar{a}la$ identification: Needleman–Wunsch (NW) Matching Score and Stroke Ratio Score.
+Each component is detailed in the following subsections.
+III-A Tabla Stroke Transcription (TST)
+This section presents our solution for TST with limited labelled data.
 We train and test deep learning models using three methods to identify the optimal transcription approach.
-Subsequently, we leverage the pre-trained model weights from the best method for transfer learning and model-agnostic meta-learning (MAML) for tabla stroke transcription.
-IV-A1 Step1: Pre-training
+Subsequently, we leverage the pre-trained model weights from the best method for MAML for TST.
+III-A1 Step1: Pre-training
 In this context, we implement three distinct transcription methods, as explained below.
 One-way Transcription Method (OTM) [6, 7]
-We adopt the method introduced by Rohit et al.
-[5, 6], expanding the stroke classes from four to ten.
-Ten distinct CNN models are employed to classify the ten stroke classes.
-Let $(X_{i},Y_{i})$ represent a single sample pair from a labeled dataset comprising $N$ samples of individual strokes.
-Here, the input $X_{i}\in\mathbb{R}^{3\times F\times T}$ to the model consists of a 3-channel chunk of 15-time frames $(T)$ extracted from a stack of 3 log-Mel spectrograms.
-The model $f_{\theta}$, parameterized by $\theta$, predicts the presence of onset for the center frame by leveraging contextual information from surrounding frames.
-It compares its output against a binary target $Y_{i}\in{0,1}$, where 1 denotes stroke onset, and 0 denotes no onset.
+We adopt the method of Rohit et al.
+[6, 7], expanding stroke classes from four to ten, with a separate CNN model for each class.
+Let $(X_{i},Y_{i})$ be a sample from a labelled dataset of $N$ strokes.
+The input $X_{i}\in\mathbb{R}^{3\times F\times T}$ is a 3-channel chunk of 15 time frames $(T)$ from log-Mel spectrograms computed using window sizes of 23ms, 46ms, and 93ms (10ms hop), stacked to combine multiple time-frequency resolutions for capturing both transient and tonal features.
+The model $f_{\theta}$ predicts the onset at the center frame by leveraging contextual information from surrounding frames, with $Y_{i}\in{\{0,1\}}$ indicating the presence or absence of a stroke onset.
 Proposed Transcription Method - 1 (PTM1)
@@
 The proposed method utilizes a CRNN for stroke transcription.
-The labeled dataset of $N$ samples is denoted as $D(X_{i},Y_{i})$, where $i=0,1,…,N$.
+The labeled dataset of $N$ samples is denoted as $D(X_{i},Y_{i})$, where $i=0,1,…,N-1$.
 Here, $X_{i}\in\mathbb{R}^{F\times T}$ is the input feature matrix (standardized log Mel-spectrogram), containing $F$ acoustic features extracted from each frame in each audio recording with $T$ time frames.
@@
 where, $\alpha\in\mathbb{R}^{+}$ is the learning rate, and $L_{wCE}$ is weighted categorical cross-entropy loss to handle the class imbalance defined as:
-$L_{wCE}=-\sum_{c=1}^{C}w_{c}y_{i}\ log(\hat{y_{i}})$ (2)
+$L_{wCE}=-\sum_{c=1}^{C}w_{c}y_{ic}\ \mathrm{log}(\hat{y_{ic}})$ (2)
 where, $C$ is the number of classes, $\hat{y_{i}}$ is the predicted output, and $y_{i}$ is the ground truth for $i^{th}$ time frame for class $c$, i.e., 0 or 1 and $w_{c}\in\mathbb{R}^{+}$ is inversely proportional to the total number of time frames in training data corresponding to each class $c$.
-Post-processing is applied to these predictions in two steps: First, if the predicted class label of $frame_{i-1}$ and $frame_{i+1}$ are the same but different from the class label of frame $frame_{i}$ then the class label of $frame_{i}$ is changed to match $frame_{i-1}$ and $frame_{i+1}$.
-In the second step, the frame where the class label changes from one class to another is identified as the onset frame for the new class.
-All onset frames in the test sample are identified, and the temporal locations of the onsets are calculated from these frames.
+After obtaining the frame-wise class predictions from the model, a post-processing step is applied to refine the output and extract onset times.
+This involves two stages.
+First, label smoothing is performed to correct short-term inconsistencies: for any frame $t\in[1,T-2]$, if the predicted class at $t$ differs from those at $t-1$ and $t+1$, but the labels at $t-1$ and $t+1$ are identical, then the label at $t$ is updated to match the surrounding frames, i.e., $Y_{c}[t]\leftarrow Y_{c}[t-1]$.
+Here, $T$ denotes the total number of time frames.
+Second, onset frames are identified as the time points where the predicted class changes between consecutive frames, i.e., $Y_{c}[t]\neq Y_{c}[t-1]$.
+These frame indices are then converted to onset times using the known frame rate, providing precise temporal localization of stroke events.
 Proposed Transcription Method - 2 (PTM2)
-[…]
-IV-A2 Step 2 - A): Model Agnostic Meta-learning (MAML)
+This method is similar to the method discussed in Section III-A1, utilizing the same CRNN model architecture and training procedure.
+However, it adopts a different target vector methodology, similar to that in [15, 16].
+The targets are represented as one-hot encoded vectors, with $Y_{i}\in\ {\{0,1\}}^{C\times T}$, where $C$ denotes the number of predefined stroke classes.
+In this method, only the onset frame of a particular class is assigned for that class, while all non-onset frames are assigned to the ‘No-stroke’ class.
+During the training phase, the model parameters are randomly initialized and updated using the stochastic gradient descent algorithm, as described by Equation (1).
+A weighted cross-entropy loss function is employed as indicated by Equation (2).
+III-A2 Step 2: Model Agnostic Meta-learning (MAML)
 The complete MAML process can be divided into meta-training and meta-testing phases.
@@
 Each task $T_{i}\sim p(T)$ is divided into a support set $T_{i}^{s}$ consisting of $s$ samples and a query set $T_{i}^{q}$ containing $q$ samples.
-The pre-trained base model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$, with parameters $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ explained in Section IV-A1, serves as the base-learner model.
-During meta-learning, we freeze the parameters of the convolutional layers and update the parameters of the recurrent and classifier layers.
+The pre-trained base model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$, with parameters $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ explained in Section III-A1, serves as the base-learner model.
+During meta-learning, we freeze the convolutional layers (pre-trained on mridangam) and update the recurrent and classifier layers.
+The convolutional layers capture low-level features shared between mridangam and tabla, preserving transferable features and reducing overfitting on low-resource tabla data.
 Thus, $\theta_{1}$ remains the same as in the pre-trained model while $\theta_{2}$ and $\theta_{3}$ become trainable parameters.
@@
 Once base learning is complete, the model $f_{[\theta_{1},\phi]}$ becomes $f_{[\theta_{1},\phi_{N}^{i}]}$.
-Subsequently, using these updated parameters $f_{[\theta_{1},\phi_{N}^{i}]}$, the query loss $L_{T_{i}^{q}}({\phi}_{N}^{i},T_{i}^{q})$ is computed on the query set $T_{i}^{q}$.
-The meta-parameters $\phi$ are then updated using the query loss.
+Subsequently, using these updated parameters $f_{[\theta_{1},\phi_{N}^{i}]}$, the loss $L_{T_{i}^{q}}(\phi_{N}^{i},T_{i}^{q})$ is computed on the query set $T_{i}^{q}$.
+The meta-parameters $\phi$ are then updated using this loss.
 The process of updating meta-parameters over the batch of tasks is called outer loop optimization (OLO) and is given by,
@@
 The model’s performance with final updated parameters $\phi_{N}^{j}$ is then evaluated on the query set $T_{j}^{q}$.
-IV-A3 Step 2 - B): Transfer learning
-We use pre-trained models from the mridangam dataset and fine-tune them with a small tabla stroke dataset to leverage their acoustic similarities.
-Transfer learning is applied to a pre-trained model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ with parameters $\theta_{1}$, $\theta_{2}$ and $\theta_{3}$, as described in Section IV-A1.
-During fine-tuning, $\theta_{1}$ remains fixed, while $\theta_{2}$ and $\theta_{3}$ are combined and updated as $\phi$.
-The parameters $\phi$ are optimized using the stochastic gradient descent algorithm Equation (1) and the weighted cross-entropy loss Equation (2).
-IV-B $T\bar{a}la$ identification
-$T\bar{a}las$ are primarily identified by their $th\bar{e}k\bar{a}$, which maintains a fixed stroke count as discussed in Section III.
+III-B $T\bar{a}la$ identification
+$T\bar{a}las$ are primarily identified by their $th\bar{e}k\bar{a}$, which maintains a fixed stroke count as discussed in Section II-C.
 We utilize these sequence and ratio properties for $t\bar{a}la$ identification, defining two $t\bar{a}la$ identification methods (scores): ‘NW Matching Score’ and ‘Stroke Ratio Score,’ based on the Needleman-Wunsch algorithm [39] and Cosine Similarity, respectively.
@@
 Figure 2: The NW Matching Score compares a transcripted sequence of $n$ strokes with a reference sequence of $m$ strokes.
-IV-B1 NW Matching Score
+III-B1 NW Matching Score
 We obtain a stroke sequence $Y=\left\langle y_{0},y_{1},...,y_{n-1}\right\rangle$ of $n$ strokes from the tabla stroke transcription module.
@@
 This approach addresses potential missing strokes in specific frames of the test sample.
-IV-B2 Stroke Ratio Score
+III-B2 Stroke Ratio Score
 We obtain the stroke count from the tabla stroke transcription module and calculate the stroke ratio for the given audio file.
@@
 where, $R$ is the reference stroke ratio vector, and $T$ is the test stroke ratio vector.
-IV-C Automatic Drum Transcription (ADT)
-This section presents our solution for ADT with minimal labelled data. We employ transfer learning and model-agnostic meta-learning (MAML) for ADT.
-IV-C1 MAML for ADT
-A similar MAML process, as explained in Section IV-A2, is adopted for ADT.
-This approach employs a CRNN model as the base learner for multi-label classification.
-The input feature matrix is a standardised log Mel-spectrogram, similar to those used in PTM1 and PTM2, with the target vector methodology following the approach described in [12, 13].
-Unlike tabla stroke transcription, where the convolutional layer parameters are frozen during meta-training and meta-testing, all parameters $\theta_{1},\theta_{2},\theta_{3}$ are trainable in ADT.
-These parameters are optimised using the weighted binary cross-entropy loss and updated via the stochastic gradient descent algorithm.
-IV-C2 Transfer Learning for ADT
-A similar transfer learning approach, as detailed in Section IV-A3, is utilized for ADT.
-The model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ is initially trained on a source domain dataset and subsequently fine-tuned on a target domain dataset.
-Model parameters are optimized using the stochastic gradient descent algorithm with the weighted binary cross-entropy loss.
+III-C Automatic Drum Transcription (ADT)
+This section presents our ADT solution under limited labelled data conditions.
+We adopt MAML for this task, following a process similar to that described in Section III-A2.
+A CRNN model serves as the base learner for multi-label classification.
+The input feature matrix is a standardized log Mel-spectrogram, consistent with those used in PTM1 and PTM2, and the target vector is constructed following the approach in [15, 16].
+Unlike in TST, where pre-trained weights are used, no relevant pre-trained model exists for Western drum kits in ADT.
+Therefore, the entire model—including convolutional, recurrent, and linear layers—is trained from scratch, with all parameters $\theta_{1},\theta_{2},\theta_{3}$ kept trainable.
+where convolutional layer parameters are frozen during meta-training and meta-testing, all parameters $\theta_{1},\theta_{2},\theta_{3}$ are trainable in ADT.
+These parameters are optimized using a weighted binary cross-entropy loss and updated via stochastic gradient descent.
+IV Datasets
+We use diverse datasets for TST and ADT, including Only Drum/Tabla (DTD), Drums with Percussion (DTP), and Complete Concert Mixes (DTM), comprising drums/tabla alongside other percussion, melodic instruments, and vocals.
+Tables I and II summarize the datasets, detailing notations, durations, number of tracks/compositions, stroke classes, and data usage for meta-training/testing.
+While some stroke or instrument classes (e.g., $Dha$, $Dhi$, snare, hi-hat) appear across datasets, they are treated as distinct during meta-learning due to differences in recording conditions, background instrumentation, and data distribution.
+Even with identical labels, contextual variations (e.g., presence of accompanying instruments or vocals) require treating them separately to ensure robust domain adaptation.
+Audio samples for meta-training and testing were randomly selected as per durations listed in Tables I and II.
+TABLE I: Meta-Learning Dataset Configuration for Tabla Stroke Transcription (TST)
+Dataset Notation Total Dur. (hr) Stroke Classes Used Meta Train-Test Data Usage & Evaluation
+Synthetic Mridangam
+Stroke Dataset
+$D_{M}$
+      2.0
+(720 excerpts)
+$Thi,Ta,Num,Tha,Dhin,Cha,$
+$Bheem,Thom,Tham,Dheem$
+Used in Pre-training stage
+Tabla Solo
+Dataset [25]
+$D_{T1}$
+      0.29
+(38 composi-
+      -tions)
+$Da,Ki,Ge,Kda,Tit,Dha,Dhe,$
+$Dhi,Re,Tin,Ta,Na,Din,Dhet,$
+$Dhin$
+Meta-Train: 12 min (0.2 hr) and 10 classes
+($Da,Ki,Ge,Kda,Tit,Dha,Dhe,Dhi,Re,Tin$)
+Meta-Test-1: 5 min (0.08 hr) adaptation on 5 new classes
+($Ta,Na,Din,Dhet,Dhin$); Evaluation on 0.21 hr
+4way-tabla-Dataset
+ismir21 [6]
+$D_{T2}$
+      1.6
+(38 tracks)
+B, D, RB, RT
+Meta-Test-2: 10 min (0.17 hr) adaptation on new classes;
+Evaluation on a 20 min (0.33 hr) separate test set by [6]
+Hindustani Music
+Rhythm Dataset [34]
+$D_{T3}$
+        5.03
+(151 excerpts)
+$Dha,Dhim,Tin,Na,Tun,Kat,$
+$Ta,Dhage,Tirikita,Dhi,Ti$
+Meta-Test-3: 10 min (0.17 hr) adaptation on new classes;
+Evaluation on 4.86 hr
+TABLE II: Meta-Learning Dataset Configuration for Automatic Drum Transcription (ADT)
+Dataset Notation ADT Task Type Total Dur. (hr) Stroke Classes Used Meta Train-Test Data Usage & Evaluation
+ADTOF [40] $D_{D4}$ DTM
+      114
+(1739 tracks)
+BD, SD, TT, HH, CY+RD Meta-Train: on full dataset (all classes seen)
+IDMT-SMT-
+Drums [41]
+$D_{D1}$ DTD
+     2.04
+(104 tracks)
+KD, SD, HH
+Meta-Test-1: 5-min (0.083hr) adaptation
+on new classes; Evaluation on 1.96 hr
+ENST-Drums
+minus-one [42]
+$D_{D2}$ DTP
+     1.01
+(64 tracks)
+BD, SD, CL, HH,
+BE, TT, RD, CY
+Meta-Test-2: 5-min (0.083 hr) adaptation
+on new classes; Evaluation on 0.927 hr
+MDB-Drums [43] $D_{D3}$ DTD, DTM
+0.35
+(23 tracks)
+KD, SD, HH, HH-O,
+TT, CY, RD, OT
+Meta-Test-3: 5-min (0.083 hr) adaptation
+on new classes; Evaluation on 0.267 hr
+IV-A Synthetic Mridangam Stroke Dataset $(D_{M})$
+The mridangam is a South Indian barrel-shaped percussion instrument with treble and bass heads, offering a wide tonal range.
+It is traditionally carved from jackfruit wood and fitted with animal skin membranes.
+We curated a dataset of 720 ten-second audio files by randomly concatenating strokes from the Mridangam Stroke Dataset [44], which contains 10 stroke classes across 6 tonic pitches, recorded at 44.1 kHz.
+The dataset includes 120 files per tonic, totalling 120 minutes of audio.
 V Experimental Setup
-This study employs simple deep CNN and CRNN models for mridangam and tabla stroke transcription and the ADT task across all experiments.
-The architectures of the CNN and CRNN models are depicted in Fig.
-3.
-Comprehensive details regarding the datasets utilized and the experimental settings for each method are outlined in the following sections.
+The experimental setup follows the same flow as shown in Fig.1 and consists of two major components: TST and ADT.
+Across all experiments, we employ simple deep CNN and CRNN models for mridangam and tabla stroke transcription, as well as for the ADT task.
+The architectures of the CNN and CRNN models are illustrated in Fig.3.
 Figure 3: General CNN [7] and CRNN model architecture for mridangam and tabla stroke transcription
-V-A Datasets
-For ADT, we use the $D_{D1}$ : IDMT-SMT-Drums [46] (DTD type), $D_{D2}$: ENST-Drums minus-one [47] (DTP type), $D_{D3}$: MDB-Drums [48] (DTD and DTM type), and $D_{D4}$ : ADTOF [49] (DTM type).
-For TST, we use the $D_{T1}$ : Tabla Solo Dataset [9], $D_{T2}$ : the 4way-tabla-ismir21-Dataset [5], and $D_{T3}$: the Hindustani Music Rhythm Dataset [43].
-$D_{T1}$ and $D_{T2}$ are tabla-only datasets, while $D_{T3}$ includes tabla along with melodic instruments and vocals, forming a complete concert dataset.
-Additionally, we curated a Synthetic Mridangam Stroke Dataset $(D_{M})$, as explained below.
-Synthetic Mridangam Stroke Dataset $(D_{M})$
-We curated a dataset of 720 audio files, each with a duration of 10 seconds, by concatenating randomly selected mridangam strokes from the existing Mridangam Stroke Dataset [50].
-This original dataset includes audio samples for ten distinct stroke labels across six different tonic pitches, recorded at a sampling rate of 44.1 kHz.
-The resulting dataset consists of 120 audio files for each tonic, amounting to a total audio duration of 120 minutes.
-TABLE I: TST : Meta-train-test dataset
-Train/Test Datasets Stroke Classes
-Meta-train
-$D_{T1}$
-Da, Ki, Ge, Kda, Tit,
-Dha, Dhe, Dhi, Re, Tin
-Meta-test-1
-$D_{T1}$ Ta, Na, Din, Dhet, Dhin
-Meta-test-2
-$D_{T2}$
-B, D, RB, RT
-Meta-test-3
-$D_{T3}$
-Dha, Dhin, Tin, Na, Tun, Kat,
-Ta, Dhage, Tirkita, Dhi, Ti
-TABLE II: ADT : Meta-train-test dataset
-Train/Test Dataset Stroke classes
-Meta-train $D_{D4}$ BD, SD, TT, HH, CY+RD
-Meta-test-1 $D_{D1}$ KD, SD, HH
-Meta-test-2 $D_{D2}$
-BD, SD, CL, HH, BE,
-TT, RD, CY
-Meta-test-3 $D_{D3}$
-KD, SD, CHH, HH, TT,
-RD, CY, OT, TMB
-V-B Tabla Stroke Transcription (TST)
-V-B1 Step - 1: Pre-training on $D_{M}$
-One-way Transcription Method
-We employ 10 CNNs for 10 classes as discussed in V-A.
+V-A Tabla Stroke Transcription (TST)
+V-A1 Step - 1: Pre-training on $D_{M}$
+One-way Transcription Method (OTM)
+We employ 10 CNNs for 10 classes as discussed in IV-A.
 Akshay et al.
@@
 We trained models on five folds and tested on one, with each fold’s training data split 90:10 for training and validation.
-This process covered all six tonics and we averaged the evaluation scores.
-The CNN processes log-Mel spectrograms from three channels with window durations of 23 ms, 46 ms, and 92 ms, all using a fixed hop size of 10 ms.
-Each channel’s input has dimensions of 128 Mel-frequency bands by 15 time frames.
-Training employs the Adam optimizer with a batch size 256 for up to 150 epochs.
-Early stopping with a patience of 10 epochs is used, saving the model with the lowest validation loss.
-During inference, a fixed-threshold peak picking algorithm [52] binarizes the output, identifying significant peaks to convert continuous values into binary decisions.
+This process covered all six tonics, and we averaged the evaluation scores.
+The CNN, trained on 3-channel log-Mel spectrogram inputs, uses the Adam optimizer with a batch size of 256 for up to 150 epochs.
+Early stopping with a patience of 10 epochs is applied, and the model with the lowest validation loss is retained.
+During inference, a fixed-threshold peak picking algorithm [46] is used to binarize the network output, identifying significant peaks and converting the continuous activations into binary decisions.
 Proposed Transcription Method - 1 (PTM1)
-The datasets used in this work vary in audio length, so we segment them into nonoverlapping 5-second chunks for feature extraction, each labeled accordingly.
+The datasets used in this work vary in audio length, so we segment them into non-overlapping 5-second chunks for feature extraction, each labelled accordingly.
 The input to the CRNN is a standardized log Mel spectrogram, computed with a 46.4 ms window size and a 10 ms hop size, utilizing 128 Mel filter banks.
@@
 Therefore, stroke onset is predominantly defined by the energetic attack-decay phase, while the sustain phase defines how the stroke ends.
-The last part of the release region contributes minimally to defining stroke onset and class.
+The last part of the release region contributes limitedly to defining stroke onset and class.
 If an equal focus is given to the complete stroke duration during transcription, it can confuse the model since this phase is similar across all stroke classes.
 Hence, we transcribe this segment separately as ‘No-stroke’ using a threshold set at 3% of the stroke’s maximum amplitude.
-Frames from onset to threshold denote the stroke class, while frames from threshold to the subsequent stroke onset are labeled ‘No-stroke.’ We employ 6-fold cross-validation using six tonics.
+Frames from onset to threshold denote the stroke class, while frames from threshold to the subsequent stroke onset are labelled ‘No-stroke.’ We employ 6-fold cross-validation using six tonics.
 Training involves a batch size of 32 over 100 epochs, with a learning rate of 0.001.
@@
 Proposed Transcription Method - 2 (PTM2)
-The CRNN model architecture and normalized log Mel spectrogram input to the CRNN are the same as explained in Section V-B1.
-The training targets include the ‘No-stroke’ class, as detailed in Section IV-A1.
+The CRNN model architecture and standardized log Mel spectrogram input to the CRNN are the same as explained in Section V-A1.
+The training targets include the ‘No-stroke’ class, as detailed in Section III-A1.
 The model is trained with a batch size of 32 for 100 epochs, using a learning rate of 0.001.
@@
 Then, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class.
-V-B2 Step 2
-Based on the pretraining results, one-way transcription performs poorly in complexity and time consumption, requiring a CNN for each stroke class.
+V-A2 Step 2
+Based on the pretraining results, OTM performs poorly in complexity and time consumption, requiring a CNN for each stroke class.
 As a result, we will discontinue this approach and instead focus on the proposed methods.
-Specifically, we will employ the CRNN model for MAML-based domain adaptation and transfer learning-based domain adaptation for tabla strokes.
+Specifically, we will employ the CRNN model for MAML-based domain adaptation for tabla strokes.
 Our methodology will be validated across three tabla datasets.
 MAML
-We meta-train our models on ten stroke classes from $D_{T1}$ and conduct meta-testing in three scenarios: evaluating performance on five stroke classes from $D_{T1}$, testing with four stroke classes from $D_{T2}$, and assessing eleven stroke classes from $D_{T3}$.
-Additionally, we performed experiments using five, six, and eleven stroke classes from $D_{T3}$ to study the impact of class number in meta-testing.
-Detailed stroke class distributions for meta-training and meta-testing are in Table I.
-We randomly selected 12 minutes of audio from $D_{T1}$ for meta-training and used 5 minutes from $D_{T1}$ , 10 minutes from $D_{T2}$ , and 10 minutes from $D_{T3}$ for meta-testing.
-Furthermore, the model adapted during meta-testing on $D_{T2}$ was directly evaluated on a separate test partition from $D_{T2}$ provided by Rohit et al.
-[5], enabling comparison with the baseline approach.
+We meta-train our models on ten stroke classes from $D_{T1}$ and evaluate in three meta-testing scenarios involving $D_{T1}$, $D_{T2}$, and $D_{T3}$.
+The corresponding stroke class distributions and audio durations for both meta-training and testing are detailed in Table I.
+All datasets come from different distributions and recording backgrounds.
+Additionally, the model adapted on $D_{T2}$ during meta-testing is directly evaluated on a separate test partition from $D_{T2}$ provided by Rohit et al.
+[6], allowing comparison with the baseline.
 In the first scenario, we must partition the existing classes into two mutually exclusive sets.
@@
 [31].
-Here, we preserve the original class label for desired strokes and classify non-target strokes as ’Others.’ As a result, there will be differences in the classes used between meta-training and meta-testing.
-Although some stroke types in $D_{T2}$ and $D_{T3}$ may appear similar, they are performed with different background instruments: $D_{T3}$ includes vocals and various instruments, while $D_{T2}$ is standalone.
-Therefore, we treat them as distinct stroke types.
-The ‘No-stroke’ class is also considered in the ‘PTM1’ and ‘PTM2’, discussed earlier in Sections IV-A1 and IV-A1, respectively.
+Here, we preserve the original class label for desired strokes and classify non-target strokes as ‘Others’.
+As a result, there will be differences in the classes used between meta-training and meta-testing.
+The ‘No-stroke’ class is also considered in the ‘PTM1’ and ‘PTM2’, discussed earlier in Sections III-A1,V-A1, and III-A1.
 The MAML approach is applied to both methods with the same experimental settings, utilizing their respective input-output pairs.
@@
 The learning rates $\alpha$ and $\beta$ are fixed at 0.001.
-Transfer Learning
-We validate the transfer learning method using the same classes and samples from $D_{T1}$, $D_{T2}$, and $D_{T3}$ as in Section V-B2.
-Additionally, a fine-tuned model on $D_{T2}$ is tested on separate test data provided by Rohit et al.
-[5] in $D_{T2}$.
-Transfer learning is applied with the same experimental settings to both methods described in Sections IV-A1 and IV-A1, using their respective input-output pairs.
-The models are trained with a batch size of 32 for 100 epochs and a learning rate of 0.001.
-V-C Automatic Drum Transcription (ADT)
-V-C1 MAML for ADT
-Meta-training is performed using five stroke classes from $D_{D4}$ while meta-testing is carried out in three distinct scenarios: (i) three stroke classes from $D_{D1}$, (ii) eight stroke classes from $D_{D2}$, and (iii) nine stroke classes from $D_{D3}$.
-The detailed distribution of stroke classes used for meta-training and meta-testing is provided in Table II.
-The complete $D_{D4}$ dataset is utilized for meta-training, and 5-minute excerpts from $D_{D1}$, $D_{D2}$, and $D_{D3}$ are used for meta-testing.
+To evaluate the effectiveness of our proposed method against existing approaches, we conducted comparative experiments using a standard supervised training setup.
+Specifically, we trained and tested the same CRNN model using the PTM1 and PTM2 paradigms, with limited training data matching the meta-adaptation setup.
+We also evaluated a transfer learning baseline, where the model was fine-tuned and tested using the same adaptation and evaluation data.
+All experiments—supervised, transfer learning, and meta-learning—were performed on identical train-test splits to ensure fair comparison.
+The models were trained with a batch size of 32 for 100 epochs using a learning rate of 0.001.
+V-A3 Baseline Models for TST Comparison
+To evaluate the effectiveness of our proposed method, we conducted comparative experiments using a standard supervised setup.
+We trained and tested the same CRNN model with PTM1 and PTM2, using limited data matching the meta-adaptation setup.
+We also evaluated transfer learning, where the model was fine-tuned and tested with the same data.
+All experiments (supervised, transfer learning, and meta-learning) used identical train-test splits for a fair comparison.
+The models were trained with a batch size of 32 for 100 epochs at a learning rate of 0.001.
+V-B Automatic Drum Transcription (ADT)
+V-B1 MAML for ADT
+Meta-training is performed using five stroke classes from $D_{D4}$, while meta-testing is conducted across three scenarios involving $D_{D1}$, $D_{D2}$, and $D_{D3}$.
+The specific stroke class distributions and audio durations for both meta-training and testing are summarized in Table II.
 For evaluation, the model trained during meta-testing is directly tested on the remaining portions of each dataset to compare its performance against the baseline approach.
-The MAML framework follows the same experimental setup as in tabla stroke transcription.
-The state-of-the-art CRNN model from [15] is also trained and evaluated on the same meta-testing data to ensure a consistent and fair performance comparison.
-V-C2 Transfer learning for ADT
-In the first step, the CRNN model is trained on $D_{D4}$ and subsequently fine-tuned using the same classes and samples from the meta-test datasets $D_{D1}$, $D_{D2}$, and $D_{D3}$.
-The fine-tuned model is then evaluated on the remaining portions of each dataset, as detailed in Section V-C1.
-The training employs a batch size of 32 over 100 epochs, with a learning rate 0.001.
-This experimental setup thoroughly assesses the model’s transfer learning capabilities across different datasets.
-VI Results and Discussions
+The MAML framework follows the same experimental setup as in the tabla stroke transcription.
+V-B2 Baseline Models for ADT Comparison
+For ADT baseline comparisons, we evaluated three approaches: a simple CRNN-based classifier [17], an NMF-based method following the setup described in [47], and a transfer learning baseline.
+NMF templates were derived from $D_{D1}$ and $D_{D2}$ and adapted across all datasets.
+To assess the impact of source separation, we applied Demucs [48] to separate mixed audio and extracted the drum track, on which all methods were re-evaluated.
+All experiments used the same set of stroke classes and train-test splits as in the MAML setup to ensure fair comparison.
 TABLE III: Step-1: Transcription Performance on $D_{M}$ (f1-score)
-Approach $Bheem$ $Cha$ $Dheem$ $Dhin$ $Num$ $Ta$ $Tha$ $Tham$ $Thi$ $Thom$ Average
-OTM 83.2 47.3 55.6 51.6 76.4 64.8 70 73.3 61.1 67.1 65.04
+Approach $Bheem$ $Cha$ $Dheem$ $Dhin$ $Num$ $Ta$ $Tha$ $Tham$ $Thi$ $Thom$ Avg.
+OTM 83.2 47.3 55.6 51.6 76.4 64.8 70.0 73.3 61.1 67.1 65.04
 PTM2 95.4 86.7 87.3 78.2 91.6 85.9 91.4 83.3 85.3 86.5 87.16
@@
   • OTM: One-way Transcription Method, PTM1:Proposed Transcription Method-1, PTM2:Proposed Transcription Method-2
+TABLE IV: TST Performance on $D_{T1}$
+          (f1-score)
+Approach $Other$ $Ta$ $Na$ $Din$ $Dhet$ $Dhin$ Avg.
+PTM2 28.2 30.7 26.3 31.4 29.5 30.3 29.40
+PTM1 34.9 36.8 33.1 38.5 35.7 37.6 36.11
+PTM2+TL 69.1 48.7 50.4 51.9 58.1 53.1 55.22
+PTM1+TL 80.1 58.9 54.9 57.7 59.9 61.5 62.17
+PTM2+ML 82.4 58.8 69.8 89.3 91.4 67.5 76.53
+PTM1+ML 87.0 71.2 76.1 89.9 92.1 71.6 81.32
+  • TL: Transfer Learning, ML: Meta Learning
+TABLE V: TST Performance on $D_{T2}$ (f1-score)
+Approach B D RB RT Avg.
+Drum-pretrained [7] 2.1 44.0 17.5 48.7 28.05
+Best set of D, RT, RB [7] 81.5 83.0 63.6 86.0 78.52
+Retrained [7] 82.7 83.6 66.9 86.6 79.95
+Retrained Rohit [6] 80.1 83.3 34.1 84.3 70.40
+PTM2 36.1 37.2 23.5 40.7 34.38
+PTM1 43.8 46.8 28.6 47.7 41.73
+PTM2 + TL 57.9 59.0 35.6 65.6 54.52
+PTM1 + TL 59.4 72.1 38.6 71.0 60.28
+PTM2 + ML 81.0 79.6 63.7 82.9 76.80
+PTM1 + ML 87.3 85.8 69.4 89.9 83.10
+TABLE VI: TST Performance on $D_{T3}$ (f1-score)
+Approach $Dha$ $Dhin$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Avg.
+PTM2 15.5 25.5 8.5 21.0 14.5 17.2 15.9 21.4 17.6 14.3 8.8 16.38
+PTM1 22.7 31.0 13.5 27.4 23.5 20.2 22.8 24.2 24.6 19.3 10.2 21.76
+PTM2+TL 19.2 36.9 9.9 34.8 16.3 21.3 26.4 39.2 27.9 22.8 23.1 25.25
+PTM1+TL 24.3 45.8 24.0 50.7 61.5 26.8 48.7 47.6 45.2 33.5 26.1 39.47
+PTM2+ML 39.4 61.2 31.3 45.5 49.0 65.2 65.6 64.1 61.2 46.9 38.1 51.59
+PTM1+ML 45.5 66.9 38 55.2 87.8 71.6 81.4 78.2 78.7 49.5 40.5 63.02
+TABLE VII: $T\bar{a}la$ Identification Performance on $D_{T3}$
+Method
+Accuracy
+Time
+RLCS0 [9] 33.6 64.2
+NW Matching Score 48.9 62.77
+Stroke Ratio Score 43.1 1.41
+VI Results and Discussions
 Performance is evaluated using the f1-score with a 50 ms collar for detecting onset positions.
@@
 VI-A Tabla stroke transcription (TST)
-Table III shows the average cross-validation results of pre-training on $D_{M}$ using the three methods from Section IV-A1.
-Predictions are made by identifying and thresholding local peaks in the network’s output activations.
-The threshold is chosen to maximize the f-score on the validation set and then applied to the test set.
-TABLE IV: Step-2: TST Performance on $D_{T1}$ (f1-score)
-Approach $Other$ $Ta$ $Na$ $Din$ $Dhet$ $Dhin$ Average
-PTM2+TL 69.1 48.7 50.4 51.9 58.1 53.1 55.22
-PTM1+TL 80.1 58.9 54.9 57.7 59.9 61.5 62.17
-PTM2+ML 82.4 58.8 69.8 89.3 91.4 67.5 76.53
-PTM1+ML 87 71.2 76.1 89.9 92.1 71.6 81.32
-TABLE V: Step-2: TST Performance on $D_{T2}$ (f1-score)
-Approach B D RB RT Average
-Drum-pretrained [6] 2.1 44 17.5 48.7 28.05
-Best set of D, RT, RB [6]
-81.5 83 63.6 86 78.52
-Retrained [6] 82.7 83.6 66.9 86.6 79.95
-Retrained Rohit [5]
-80.1 83.3 34.1 84.3 70.4
-PTM2 + TL 57.9 59 35.6 65.6 54.52
-PTM1 + TL 59.4 72.1 38.6 71 60.28
-PTM2 + ML 81 79.6 63.7 82.9 76.8
-PTM1 + ML 87.3 85.8 69.4 89.9 83.1
-TABLE VI: Step-2: TST Performance on $D_{T3}$ (11 stroke classes) (f1-score)
-Approach $Dha$ $Dhin$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Average
-PTM2+TL 19.2 36.9 9.9 34.8 16.3 21.3 26.4 39.2 27.9 22.8 23.1 25.25
-PTM1+TL 24.3 45.8 24 50.7 61.5 26.8 48.7 47.6 45.2 33.5 26.1 39.47
-PTM2+ML 39.4 61.2 31.3 45.5 49 65.2 65.6 64.1 61.2 46.9 38.1 51.59
-PTM1+ML 45.5 66.9 38 55.2 87.8 71.6 81.4 78.2 78.7 49.5 40.5 63.02
-  • PTM1:Proposed Transcription Method-1, PTM2:Proposed Transcription Method-2, TL:Transfer Learning, ML: Meta Learning
-TABLE VII: Step-2: TST Performance on $D_{T3}$ (with 5 stroke classes) (f1-score)
-Approach $Other1$ $Dhin$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Average
-PTM2+TL 21.6 38.3 33.8 26.9 15.4 18.2 25.70
-PTM1+TL 35.8 43.4 45.7 48.1 34.6 22.3 38.31
-PTM2+ML 50.4 61.9 57.1 59.8 45.6 37.9 52.11
-PTM1+ML 61.4 68.8 75.7 79.7 47.6 43.9 62.85
-TABLE VIII: Step-2: TST Performance on $D_{T3}$ (with 6 stroke classes) (f1-score)
-Approach $Other2$ $Dha$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ Average
-PTM2+TL 31.5 17.2 13 37.3 19.4 24.6 28.9 24.55
-PTM1+TL 42.3 19.5 22.6 47.7 59.1 25.8 50.3 38.19
-PTM2+ML 54.1 40.1 29.7 47.7 48.9 66.3 67.4 50.6
-PTM1+ML 60.1 46.2 36.7 56.8 85.9 72.3 82.4 62.91
-TABLE IX: $T\bar{a}la$ identification performance on $D_{T3}$
-Method Accuracy (%) Time (ms)
-RLCS0[9] 33.6 64.2
-NW Matching Score 48.9 62.77
-Stroke Ratio Score 43.1 1.41
-TABLE X: ADT performance on DTD - $D_{D1}$
-(f1-score)
-Approach KD SD HH Average
-SOTA CRNN [15] 42.1 65.3 58.8 55.4
-CRNN + TL 54.3 68.7 63.9 62.3
-CRNN + ML 73.9 84.2 82.1 80.1
-TABLE XI: ADT performance on DTP - $D_{D2}$ (f1-score)
-Approach BD SD CL HH BE TT RD CY Average
-SOTA
-CRNN [15]
-54.4 53.1 41.5 56.5 34.7 42.6 47.5 39.1 46.2
-CRNN + TL 68.7 65.3 52.1 64.4 48.3 61.5 58.4 53.5 59.0
-CRNN + ML 83.5 82.8 72.1 85.3 68.5 77.2 78 73.9 77.6
-TABLE XII: ADT performance on DTD - $D_{D3}$ (f1-score)
-Approach KD SD CHH HH TT RD CY OT TMB Average
-SOTA CRNN [15] 44.8 46.3 38.7 33.1 27.4 39.3 32.6 24.5 14.4 33.4
-CRNN + TL 54.3 70 63.5 60.2 57.5 62.7 54.1 44.8 33.4 55.6
-CRNN + ML 73.1 91.6 84.4 82.3 70.9 81.6 74.5 67.3 59.5 76.1
-TABLE XIII: ADT performance on DTM - $D_{D3}$ (f1-score)
-Approach KD SD CHH HH TT RD CY OT TMB Average
-SOTA CRNN [15] 41.7 38.2 29.8 30.3 24.5 36.2 29.4 18.6 8.6 28.6
-CRNN + TL 52.7 66.1 61.7 57.8 53 60.3 50.1 40.9 31.8 52.7
-CRNN + ML 71.4 87.8 80.1 79.8 68.6 78.5 64.9 61.3 54.2 71.8
-TABLE XIV: ADT results Comparison
-(Proposed vs. [36])
-Approach % f1-score
-Trained on SLAKH [36] 50.0
-Trained on STAR [36] 62.0
-Dynamic FSL [36] 70.0
-CRNN + ML 71.8
-The results demonstrate that the proposed methods outperform the existing OTM.
-The OTM, utilizing a CNN, only considers the spatial dimension of the Mel spectrogram.
-In contrast, the proposed CRNN method integrates both spatial and temporal dimensions, capturing the attack-decay-sustain-release (ADSR) temporal model of strokes more effectively.
-PTM1, which utilizes all stroke frames, outperforms PTM2, which relies on a single frame.
-PTM2’s reliance on a single frame may lead to missed stroke detections if the initial frame is undetected, whereas PTM1 remains robust by considering subsequent frames.
-Table V shows meta-learning and transfer learning results using PTM1 and PTM2 on test data $D_{T1}$ from the same domain, demonstrating PTM1 with meta-learning outperforms PTM2 and others.
-Similarly, Table V presents results on test data $D_{T2}$ from a similar domain, with PTM1 again outperforming PTM2.
-Tables VI to VIII display results on vastly different domain test data $D_{T3}$, confirming PTM1 with meta-learning consistently outperforms PTM2 across different class variations.
-Tables V-VIII demonstrate that the meta-learning consistently outperforms transfer learning.
-MAML learns an initialization for rapid task adaptation through meta-learning, while transfer learning directly uses pre-trained weights, which may not optimize performance for the target task.
-Specifically, Table V shows PTM1 with meta-learning outperforming four OTM settings using transfer learning and data augmentations, attributed to CNNs’ limited ability to capture temporal relationships.
-Tables VI-VIII indicate performance degradation on $D_{T3}$ due to frequent tabla stroke overlap with vocals and other instruments in concert data, complicating stroke onset detection.
-The number of stroke classes tested shows minimal impact on performance across these analyses.
+Table IV presents the average cross-validation results of pre-training on $D_{M}$ using the methods from Section III-A1.
+Predictions are obtained by thresholding local peaks in network activations, with thresholds tuned on the validation set.
+The proposed methods outperform OTM, which captures only frequency-domain features.
+In contrast, the CRNN leverages both spectral and temporal information, modelling the ADS characteristics of strokes.
+PTM1, which uses all frames within a stroke, outperforms PTM2, which relies solely on onset frames—likely affected by annotation inconsistencies.
+Leveraging temporal context, PTM1 offers improved robustness.
+Additionally, CRNN models perform well with single-channel input, indicating that OTM’s 3-channel configuration is not essential.
+Tables IVVII show results for PTM1 and PTM2 under meta-learning across test sets with varying data distributions.
+PTM1 consistently outperforms PTM2 and other baselines, demonstrating the advantage of MAML’s rapid task adaptation over fixed representations.
+Table V highlights PTM1 outperforming four OTM settings, emphasizing CNN’s limitations in modeling temporal dependencies.
+Performance drop on $D_{T3}$ (Table VII) is due to stroke overlap with vocals and instruments, hindering onset detection.
+The comparison with NMF-based and source separation methods is excluded for TST due to several limitations.
+Unlike ADT, TST lacks isolated instrument recordings, and using templates from the test dataset would create an unfair comparison.
+Additionally, no publicly available source separation model is designed for Indian classical music, and applying existing models like Demucs [48] leads to artifacts and domain mismatch, making them unsuitable for accurate stroke-level transcription.
 VI-B $T\bar{a}la$ identification
@@
 VI-C Automatic drum transcription (ADT)
-The performance of the proposed Transfer Learning and MAML-based approaches for ADT tasks is summarized in Tables X-XIII.
-These tables present results on datasets representing all ADT task types--DTD, DTP, and DTM--specifically $D_{DD1}$, $D_{DD2}$, and $D_{DD3}$.
-The comparison is made against the state-of-the-art CRNN model [15].
-Additionally, Table XIV highlights the comparative performance of the proposed MAML approach with existing methods on $D_{D3}$ (DTM).
-The results demonstrate that the proposed MAML-based method consistently outperformed existing approaches utilizing minimal target data across all ADT task types.
-This indicates the effectiveness of the proposed approach in scenarios with limited data availability.
+Tables VIIIXI present the performance of the proposed MAML-based approach across all ADT task types—DTD, DTP, and DTM—compared with baselines including the state-of-the-art CRNN model [17], an NMF-based method [47], and transfer learning.
+For the DTM dataset, source separation [48] is applied, and results are reported accordingly.
+Table XII further compares MAML on $D_{D3}$ (DTM) against existing methods, using a 30ms collar for fair evaluation.
+Among the baselines, the NMF-based method performs better than the supervised CRNN in low-data scenarios, showing greater robustness under limited supervision.
+However, its performance degrades on the DTM dataset due to interference from other instruments and vocals, which hampers reliable template construction in the absence of isolated tracks.
+Overall, the MAML-based approach consistently outperforms all baselines across ADT tasks, demonstrating strong generalization and robustness in low-resource and real-world conditions.
+TABLE VIII: ADT Performance on DTD - $D_{D1}$ (f1-score)
+Approach KD SD HH Avg.
+SOTA-CRNN [17] 42.1 65.3 58.8 55.4
+PF-NMF [47] 68.4 72.5 73.0 71.3
+CRNN + TL 54.3 68.7 63.9 62.3
+CRNN + ML 73.9 84.2 82.1 80.1
+TABLE IX: ADT Performance on DTP - $D_{D2}$ (f1-score)
+Approach BD SD CL HH BE TT RD CY Avg.
+SOTA-CRNN [17] 54.4 53.1 41.5 56.5 34.7 42.6 47.5 39.1 46.2
+PF-NMF [47] 77.4 66.9 69.8 74.5 63.2 71.8 69.1 68.3 70.1
+CRNN + TL 68.7 65.3 52.1 64.4 48.3 61.5 58.4 53.5 59.0
+CRNN + ML 83.5 82.8 72.1 85.3 68.5 77.2 78.0 73.9 77.6
+TABLE X: ADT Performance on DTD - $D_{D3}$ (f1-score)
+Approach KD SD HH HH-O TT CY RD OT Avg.
+SOTA-CRNN [17] 43.0 44.5 37.2 31.8 26.3 37.7 31.3 23.5 34.4
+PF-NMF [47] 59.2 65.1 71.8 63.3 68.2 66.1 63.5 32.2 61.2
+CRNN + TL 53.0 68.3 62.0 58.8 56.2 61.2 52.8 43.7 57.0
+CRNN + ML 72.3 90.5 83.4 81.3 70.1 80.6 73.6 66.5 77.3
+TABLE XI: ADT Performance on DTM - $D_{D3}$ (f1-score)
+Approach KD SD HH HH-O TT CY RD OT Avg.
+SOTA-CRNN [17] 40.4 37.0 28.9 29.3 23.7 35.1 28.5 18.0 30.1
+PF-NMF [47] 41.3 39.7 35.6 33.8 30.4 36.9 32.6 24.9 34.4
+CRNN + TL 51.3 64.3 60.0 56.2 51.5 58.6 48.7 39.8 53.8
+CRNN + ML 70.0 86.1 78.5 78.2 67.3 76.9 63.6 60.1 72.6
+SS + SOTA-CRNN 44.8 40.1 33.0 31.5 26.8 36.3 29.4 21.1 32.9
+SS + PF-NMF 54.0 68.5 64.5 61.0 56.2 64.8 60.1 52.4 60.2
+SS + CRNN + TL 53.1 67.7 62.3 57.5 53.4 60.2 51.6 42.9 56.1
+SS + CRNN + ML 72.8 88.5 80.4 79.2 70.1 78.6 66.3 61.7 74.7
+  • Note: Results use a 30ms collar; SS: Source Separation using Demucs [48]
+TABLE XII: Comparison of ADT Performance on DTM - $D_{D3}$ with Existing Methods
+Approach % f1-score
+Proto-BL8-OFF [32] 67.0
+Dense-BL8-OFF [32] 67.0
+Proto-BL21-OFF [32] 69.0
+Dense-BL21-OFF [32] 70.0
+SOTA-CRNN [17] 30.1
+SS + SOTA-CRNN 32.9
+PF-NMF [47] 34.4
+SS + PF-NMF 60.2
+CRNN + TL 53.8
+SS + CRNN + TL 56.1
+CRNN + ML 72.6
+SS + CRNN + ML 74.7
+  • Note: Results use a 30ms collar and the same classes for fair comparison
+                      with [32]; SS: Source Separation using Demucs [48]
 VII Conclusion
-This paper presents a novel approach leveraging Model-Agnostic Meta-Learning (MAML) for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT), addressing the challenge of limited annotated datasets.
-The proposed method effectively transcribes tabla strokes and drum sounds across diverse datasets, including scenarios with drum or tabla strokes alone, percussion instruments with melodic accompaniments, and the singer’s voice, ensuring comprehensive validation and demonstrating robustness and adaptability.
-Following TST, two new $t\bar{a}la$ identification methods are introduced, contributing to rhythmic analysis in Hindustani music.
-Experimental results highlight the superiority of the MAML-based approach over existing methods, consistently achieving better performance with minimal labelled data.
-Future work will focus on improving transcription accuracy for concert datasets with minimal labeled data, leveraging advanced music source separation techniques and active learning strategies.
+This paper proposes a novel Model-Agnostic Meta-Learning (MAML) based approach for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT), addressing challenges of limited annotated data and label heterogeneity.
+The method demonstrates effective transcription across diverse scenarios, including solo percussion, percussion with melodic accompaniments, and vocals, ensuring comprehensive validation and highlighting its robustness and adaptability.
+Following TST, two novel $t\bar{a}la$ identification methods are introduced, advancing rhythmic analysis in Hindustani music.
+Experimental results show that the MAML-based approach consistently outperforms existing methods, even with limited labelled data.
+Future work will focus on enhancing transcription accuracy for Indian music concert datasets with limited labelled data by leveraging advanced source separation techniques tailored to Indian music.
+We also aim to curate a larger and more diverse collection of polyphonic recordings to support training and evaluation in realistic conditions.